### v2

# Renormalizable Spectral-Shell Dynamics as the Origin of Neural Scaling Laws

Abstract
Neural scaling laws and double-descent phenomena suggest that deep-network training obeys a simple macroscopic structure despite highly nonlinear optimization dynamics. We derive such structure directly from gradient descent in function space. For mean-squared error loss, the training error evolves as $\dot{e}_{t}=-M(t)e_{t}$ with $M(t)=J_{\theta(t)}J_{\theta(t)}^{\!*}$, a time-dependent self-adjoint operator induced by the network Jacobian. Using Kato perturbation theory, we obtain an exact system of coupled modewise ODEs in the instantaneous eigenbasis of $M(t)$.
To extract macroscopic behavior, we introduce a logarithmic spectral-shell coarse-graining and track quadratic error energy across shells. Microscopic interactions within each shell cancel identically at the energy level, so shell energies evolve only through dissipation and external inter-shell interactions. We formalize this via a renormalizable shell-dynamics assumption, under which cumulative microscopic effects reduce to a controlled net flux across shell boundaries.
Assuming an effective power-law spectral transport in a relevant resolution range, the shell dynamics admits a self-similar solution with a moving resolution frontier and explicit scaling exponents. This framework explains neural scaling laws and double descent, and unifies lazy (NTK-like) training and feature learning as two limits of the same spectral-shell dynamics.
1 Introduction
Modern deep neural networks are trained by strongly nonlinear and high-dimensional optimization procedures. Nevertheless, across architectures, datasets, and training recipes, their performance exhibits remarkably simple and robust regularities. Most prominently, empirical studies have shown that training and test losses obey approximate power laws in model size, dataset size, or total compute over wide dynamic ranges (hestness2017deep; kaplan2020scaling; henighan2020scaling; hoffmann2022training; hernandez2021scaling; wei2022emergent). Related law-like behaviors have been observed in precision scaling, pruning, sparsification, and model densification (kumar2024scaling; sorscher2022beyond; rosenfeld2021predictability; blalock2020state; Xiao2024DensingLaw). These phenomena strongly suggest the existence of an underlying macroscopic structure governing how error is redistributed and dissipated during training.
Limitations of existing theories.
A classical theoretical explanation is provided by the neural tangent kernel (NTK) framework (jacot2018ntk; lee2019wide; bietti2021inductive), which models training as a linear dynamics in function space with a fixed kernel. While this perspective explains certain spectral biases and convergence rates, it relies on a frozen feature map and breaks down once representation learning becomes significant. Recent work has therefore emphasized feature-learning dynamics, using mean-field, tensor-program, or dynamical systems approaches (yang2021tensor; canatar2022kernel; bordelon2024dynamical; bordelon2023feature; wang2023spectralevolution). These models successfully capture richer behavior beyond the lazy regime, but typically treat NTK-like and feature-learning regimes via distinct approximations and do not provide a single operator-level equation that interpolates continuously between them.
A spectral-shell perspective.
In this work, we take a function-space and operator-theoretic viewpoint. For mean-squared error loss, the training error $e_{t}$ evolves according to the exact linear equation
$\dot{e}_{t}=-M(t)e_{t},\qquad M(t)=J_{\theta(t)}J_{\theta(t)}^{\!*},$
where $M(t)$ is a time-dependent, self-adjoint operator induced by the network Jacobian. Rather than assuming fixed features or a prescribed kernel, we analyze the training dynamics directly through the evolving spectral structure of $M(t)$.
Using Kato perturbation theory, we derive an exact system of coupled modewise ODEs in the instantaneous eigenbasis of $M(t)$. The resulting dynamics are fully rigorous but highly nonlocal across modes. To extract macroscopic behavior, we introduce a logarithmic spectral-shell coarse-graining and track the evolution of quadratic error energy within each shell. A key structural fact is that microscopic interactions within a shell, arising from eigenbasis drift, are strictly antisymmetric and cancel identically at the level of quadratic energy. Consequently, shell energies evolve only through dissipation and external interaction effects—namely, energy exchange with other shells.
Renormalizable shell dynamics.
This observation leads to an exact shell-level energy balance law. To close it at macroscopic scales, we introduce a renormalizable shell-dynamics assumption: after coarse-graining, the cumulative effect of microscopic degrees of freedom can be absorbed into a controlled net flux across shell boundaries, with subleading corrections becoming negligible. This notion of renormalizability is inspired by analogous constructions in statistical physics and condensed-matter theory, and it does not require a continuum limit nor impose any a priori direction of spectral transport.
When spectral shells are sufficiently dense over a dynamically relevant range, the shell dynamics admits an continuum approximation in the spectral variable $\lambda$, yielding an effective transport–dissipation equation. We emphasize that this PDE is a convenient representation of shell dynamics, not its fundamental starting point. Throughout the paper, renormalizability and effective power-law transport are treated as macroscopic assumptions, not as consequences of weak coupling or locality.
Effective transport and scaling laws.
We further assume that, in the relevant resolution range, the renormalized spectral flux can be summarized by an effective power-law transport form. Under this single coarse-grained assumption, the shell dynamics admits a self-similar solution with a moving resolution frontier, geometric amplitude growth, and power-law dissipation. This structure yields explicit scaling-law exponents and provides a unified explanation for neural scaling laws and double-descent phenomena.
Within this framework, lazy (NTK-like) training and feature learning arise as two limits of the same spectral-shell dynamics. When the effective transport vanishes, the dynamics reduce to pure dissipation with fixed features. When transport is active, spectral mass flows across resolutions, inducing representation learning. Thus, both regimes—and the continuum between them—are unified by a single operator-level shell dynamics governing the redistribution of error during training.
2 Preliminaries: Function-Space Dynamics and Error Evolution
We consider supervised learning with mean-squared error (MSE) loss in the Hilbert space $H=L^{2}(p)$ induced by the data distribution $p(x)$. All functions are identified up to sets of measure zero, and the inner product is defined by
$\langle f,g\rangle:=\mathbb{E}_{x\sim p}\,[\,f(x)\,g(x)\,],\qquad\|f\|_{2}^{2}=\langle f,f\rangle.$
2.1 Neural networks as functions in $L^{2}(p)$
A neural network with parameters $\theta\in\mathbb{R}^{N}$ represents a time-dependent function $f_{\theta(t)}\in H$. Given a target function $f^{*}\in H$, we define the error at time $t$ as
$e_{t}:=f^{*}-f_{\theta(t)}.$
Throughout this work, we emphasize that no structural assumptions such as fixed features, invariant kernels, or predetermined mode decompositions are imposed: the representation $f_{\theta(t)}$ is allowed to evolve freely in $H$, and all structure in the dynamics arises directly from gradient descent.
2.2 Gradient descent induces a linear evolution in function space
For MSE loss,
$\mathcal{L}(\theta)=\frac{1}{2}\,\|f^{*}-f_{\theta}\|_{2}^{2}=\frac{1}{2}\,\|e\|_{2}^{2},$
the gradient flow in parameter space is
$\dot{\theta}(t)=-\nabla_{\theta}\mathcal{L}(\theta(t)).$
By the chain rule,
$\dot{e}_{t}=-J_{\theta(t)}\,\dot{\theta}(t),$
where $J_{\theta}\colon\mathbb{R}^{N}\to H$ denotes the Jacobian operator $(J_{\theta}v)(x)=\nabla_{\theta}f_{\theta}(x)\cdot v$. Substituting the gradient flow equation gives
$\dot{e}_{t}=-J_{\theta(t)}\,J_{\theta(t)}^{\!*}\,e_{t}.$
Thus the error evolves according to a linear operator in function space:
$\boxed{\dot{e}_{t}=-M(t)\,e_{t},\qquad M(t):=J_{\theta(t)}\,J_{\theta(t)}^{\!*}.}$
The operator $M(t)$ is self-adjoint and positive semidefinite. It is the only operator that governs the evolution of the error in our analysis; its time dependence reflects the evolving representation of the network under gradient descent.
In this section we work in the continuous-time gradient flow limit,
$\dot{\theta}(t)=-\nabla_{\theta}\mathcal{L}(\theta(t)),$
which corresponds to taking the learning rate to be infinitesimal. A discrete-time update with (possibly time-varying) step sizes $\eta_{k}$,
$\theta_{k+1}=\theta_{k}-\eta_{k}\nabla_{\theta}\mathcal{L}(\theta_{k}),$
can be viewed as a time reparameterization of this flow, where the effective training time is proportional to the accumulated step size $\sum_{k}\eta_{k}$. Thus, learning rate schedules do not change the form of the operator dynamics $\dot{e}_{t}=-M(t)e_{t}$; instead, they induce a non-uniform rescaling of time that will later be absorbed into the effective time variable $\tau(t)$ in our spectral analysis.
2.3 Spectral decomposition of $M(t)$
For each fixed $t$, the operator $M(t)$ acts on a finite-dimensional subspace of $H$ determined by the network’s Jacobian, and therefore admits a discrete spectral decomposition. We write
$M(t)\,\varphi_{u}(t)=\lambda_{u}(t)\,\varphi_{u}(t),\qquad u\in\mathcal{U}(t),$
where $\mathcal{U}(t)$ is a finite or countable index set, the eigenvalues $\lambda_{u}(t)\geq 0$, and the eigenfunctions $\{\varphi_{u}(t)\}_{u\in\mathcal{U}(t)}$ form an orthonormal family in $H$. Expanding the error in this moving eigenbasis gives
$e_{t}=\sum_{u\in\mathcal{U}(t)}g_{u}(t)\,\varphi_{u}(t).$
The amplitudes $g_{u}(t)$ encode how much error lies along each instantaneous mode of $M(t)$, and the time dependence of both $\lambda_{u}(t)$ and $\varphi_{u}(t)$ reflects the evolution of the network’s representation.
2.4 Why a spectral formulation?
Although $M(t)$ generally has low rank compared to the ambient dimension of $H$, its eigenstructure provides a natural lens on the error dynamics. The eigenvalues $\lambda_{u}(t)$ quantify the rate at which error aligned with mode $u$ is dissipated, while the evolution of $\varphi_{u}(t)$ captures the “feature learning” aspect of training. In Section 3.1, we show that the amplitudes $g_{u}(t)$ satisfy a rigorous, coupled system of mode ODEs derived from Kato’s perturbation theory. In Section 3, we then coarse-grain these discrete modes to obtain a continuous spectral PDE that describes the statistical redistribution of error across many nearby modes.
3 From Modewise ODEs to Coarse-Grained Spectral Dynamics
In this section we derive the fundamental coarse-grained spectral structure governing the dynamics of the error under gradient descent. Starting from the exact function-space evolution $\dot{e}_{t}=-M(t)e_{t}$, we first express the dynamics in the instantaneous eigenbasis of $M(t)$, obtaining a fully rigorous system of coupled modewise ODEs.
Rather than attempting to track individual modes, which is intractable at macroscopic scales, we then introduce a logarithmic spectral-shell coarse-graining and study the evolution of the quadratic error energy carried by each shell. This leads to an exact shell-level balance law, in which shell-internal coupling cancels identically and all nontrivial interactions appear as inter-shell energy fluxes.
When the shell spacing is sufficiently fine over a dynamically relevant range, this shell dynamics admits an continuum approximation in the spectral variable $\lambda$, yielding a transport–dissipation PDE. However, the shell-level formulation itself is primary and does not rely on any continuum limit. In particular, we never assume that the spectrum becomes continuous or that a limit as the shell spacing vanishes exists; all continuum expressions should be read as local approximations to discrete shell differences.
3.1 Exact Mode ODEs in a Drifting Eigenbasis
For each time $t$, the operator $M(t)=J_{\theta(t)}J_{\theta(t)}^{\!*}$ is self-adjoint and positive semidefinite, acting on a finite-dimensional subspace of $H=L^{2}(p)$. Thus its spectrum is discrete. Let
$M(t)\varphi_{u}(t)=\lambda_{u}(t)\,\varphi_{u}(t),\qquad u\in\mathcal{U}(t),$
denote an orthonormal eigenbasis of $M(t)$. Expanding the error in this moving basis gives
$e_{t}=\sum_{u}g_{u}(t)\,\varphi_{u}(t),\qquad g_{u}(t)=\langle e_{t},\varphi_{u}(t)\rangle.$
Differentiating in time yields
$\partial_{t}g_{u}(t)=\langle\partial_{t}e_{t},\,\varphi_{u}(t)\rangle+\big\langle e_{t},\,\partial_{t}\varphi_{u}(t)\big\rangle.$
Using $\dot{e}_{t}=-M(t)e_{t}$, the first term becomes
$\langle-M(t)e_{t},\,\varphi_{u}\rangle=-\lambda_{u}(t)g_{u}(t).$
The second term encodes the rotation of the eigenbasis. From Kato’s perturbation theory (kato2012short; Zwiebach_Adiabatic) for differential self-adjoint operators, the evolution of the eigenfunctions satisfies
$\langle\partial_{t}\varphi_{v}(t),\,\varphi_{u}(t)\rangle=\begin{cases}\displaystyle\frac{\langle\varphi_{u}(t),\,\dot{M}(t)\,\varphi_{v}(t)\rangle}{\lambda_{v}(t)-\lambda_{u}(t)},&v\neq u,\\[10.00002pt] 0,&v=u,\end{cases}$
where $\dot{M}(t)$ is the operator derivative of $M(t)$. Therefore
$\big\langle e_{t},\,\partial_{t}\varphi_{u}(t)\big\rangle=\sum_{v\neq u}g_{v}(t)\,\frac{\langle\varphi_{u}(t),\,\dot{M}(t)\,\varphi_{v}(t)\rangle}{\lambda_{v}(t)-\lambda_{u}(t)}.$
Combining both contributions, we obtain the exact coupled mode ODE:
$\boxed{\partial_{t}g_{u}(t)+\sum_{v\neq u}g_{v}(t)\,\Omega_{v\to u}(t)=-\lambda_{u}(t)\,g_{u}(t),}$
where the coupling coefficients are
$\Omega_{v\to u}(t)=\frac{\langle\varphi_{u}(t),\,\dot{M}(t)\,\varphi_{v}(t)\rangle}{\lambda_{v}(t)-\lambda_{u}(t)}.$
This ODE system is fully rigorous and contains all aspects of the dynamics: local dissipation $\lambda_{u}g_{u}$, nonlocal mode coupling $\sum_{v\neq u}g_{v}\,\Omega_{v\to u}$, and the drift of the feature basis through the time dependence of the eigenfunctions $\varphi_{u}(t)$.
3.2 Logarithmic Spectral Shells and Shell Energies
The exact modewise dynamics derived in Section 3.1 are fully rigorous but too fine-grained for macroscopic analysis. To expose the coarse structure, we group modes into logarithmic spectral shells and track the quadratic error energy carried by each shell.
Fix a ratio $q>1$ and define a logarithmic partition of the positive spectrum:
$\lambda_{\alpha}:=\lambda_{0}\,q^{\alpha},\qquad S_{\alpha}:=\bigl\{\,u:\ \lambda_{u}(t)\in[\lambda_{\alpha},\lambda_{\alpha+1})\,\bigr\},\qquad\alpha\in\mathbb{Z}.$
We also define cumulative shells $S_{\leq\alpha}:=\bigcup_{\gamma\leq\alpha}S_{\gamma}$.
For each shell, we define the shell quadratic energy
$E_{\alpha}(t):=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}(t)^{2},\qquad E_{\leq\alpha}(t):=\sum_{\gamma\leq\alpha}E_{\gamma}(t).$
This choice is canonical: the full function-space loss is $\mathcal{L}(t)=\frac{1}{2}\|e_{t}\|_{2}^{2}=\frac{1}{2}\sum_{u}g_{u}(t)^{2}$, so $\{E_{\alpha}\}$ provides an exact energy bookkeeping across spectral shells.
3.3 Exact Cancellation of Shell-Internal Coupling for Quadratic Energy
Recall the exact mode ODE
$\partial_{t}g_{u}(t)+\sum_{v\neq u}g_{v}(t)\,\Omega_{v\to u}(t)=-\lambda_{u}(t)\,g_{u}(t),\qquad\Omega_{v\to u}(t)=\frac{\langle\varphi_{u}(t),\dot{M}(t)\varphi_{v}(t)\rangle}{\lambda_{v}(t)-\lambda_{u}(t)}.$
Since $\dot{M}(t)$ is self-adjoint and the denominator is antisymmetric, we have the strict antisymmetry
$\Omega_{v\to u}(t)=-\Omega_{u\to v}(t),\qquad u\neq v.$
Proposition 1 (Shell-internal cancellation (quadratic energy)).
For any spectral shell $S_{\alpha}$, the shell-internal coupling does not change the quadratic shell energy:
$\sum_{u\in S_{\alpha}}\sum_{v\in S_{\alpha},\ v\neq u}g_{v}(t)\,g_{u}(t)\,\Omega_{v\to u}(t)=0.$
Proof.
Pair terms $(u,v)$ and $(v,u)$ inside the double sum. Using $\Omega_{v\to u}=-\Omega_{u\to v}$, we get
$g_{v}g_{u}\Omega_{v\to u}+g_{u}g_{v}\Omega_{u\to v}=g_{u}g_{v}\bigl(\Omega_{v\to u}+\Omega_{u\to v}\bigr)=0.$
Summing over all unordered pairs in $S_{\alpha}$ yields the claim. ∎
Remark 1 (Interpretation).
Proposition 1 formalizes the key point: shell-internal coupling is a pure redistribution (rotation) of error among modes within the shell. It can change the individual $g_{u}$’s, but it cannot change $\sum_{u\in S_{\alpha}}g_{u}^{2}$. Therefore, any change of the shell energy $E_{\alpha}(t)$ must come from (i) local dissipation and (ii) energy exchange with other shells.
3.4 Exact Shell Balance Law for Quadratic Energy
Differentiate $E_{\alpha}(t)=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}^{2}$:
$\frac{d}{dt}E_{\alpha}(t)=\sum_{u\in S_{\alpha}}g_{u}(t)\,\partial_{t}g_{u}(t).$
Substituting the mode ODE yields
$\frac{d}{dt}E_{\alpha}(t)$ $=-\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2}-\sum_{u\in S_{\alpha}}\sum_{v\neq u}g_{v}(t)\,g_{u}(t)\,\Omega_{v\to u}(t).$ (1)
Split the coupling sum into $v\in S_{\alpha}$ and $v\notin S_{\alpha}$. By Proposition 1, the $v\in S_{\alpha}$ contribution cancels. Hence we obtain the exact shell energy balance:
$\boxed{\frac{d}{dt}E_{\alpha}(t)=-\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2}-\sum_{\beta\neq\alpha}\ \sum_{\begin{subarray}{c}u\in S_{\alpha}\\ v\in S_{\beta}\end{subarray}}g_{v}(t)\,g_{u}(t)\,\Omega_{v\to u}(t).}$ (2)
The first term is pure shell dissipation. The second term is pure inter-shell exchange.
3.5 Inter-Shell Fluxes and Discrete Conservation Structure
Define the inter-shell quadratic-energy flux from $S_{\beta}$ to $S_{\alpha}$ by
$\mathcal{F}_{\beta\to\alpha}(t):=-\sum_{\begin{subarray}{c}u\in S_{\alpha}\\ v\in S_{\beta}\end{subarray}}g_{v}(t)\,g_{u}(t)\,\Omega_{v\to u}(t).$
Then equation (2) becomes
$\boxed{\frac{d}{dt}E_{\alpha}(t)=-\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2}+\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t).}$ (3)
Moreover, antisymmetry implies a strict action–reaction identity:
$\mathcal{F}_{\beta\to\alpha}(t)=-\mathcal{F}_{\alpha\to\beta}(t).$
Therefore, coupling conserves total quadratic energy across shells:
$\sum_{\alpha}\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t)=0,$
and the only mechanism that decreases $\sum_{\alpha}E_{\alpha}(t)=\frac{1}{2}\|e_{t}\|_{2}^{2}$ is the dissipation term $\sum_{u}\lambda_{u}g_{u}^{2}$.
3.6 From Exact Shell Balance to Renormalized Dynamics
The balance law (3) is an exact conservation-type bookkeeping equation across logarithmic shells, requiring no continuum limit and no PDE interpretation. However, the fluxes $\mathcal{F}_{\beta\to\alpha}(t)$ remain nonlocal and depend on microscopic details of the evolving operator.
In the remainder of the paper, we do not attempt to characterize $\mathcal{F}_{\beta\to\alpha}$ microscopically. Instead, Section 4 introduces a renormalizable shell-dynamics assumption that closes the cumulative effect of inter-shell fluxes at the shell level. All subsequent results rely exclusively on this shell-level interface.
We emphasize that all subsequent results in this paper rely exclusively on this shell-level energy balance and the renormalizable flux interface introduced in Section 4. The continuum PDE description is used only as a convenient approximation when shell resolution permits.
4 Renormalizable Shell Dynamics and Effective Power-Law Transport
Section 3 provides an exact modewise ODE system and an effective spectral PDE description. In this section we introduce an explicitly coarse-grained interface for macroscopic analysis: logarithmic spectral shells, a renormalizable flux bookkeeping condition across shells, and a single effective transport assumption used in the remainder of the paper. We emphasize that this section is not a microscopic characterization of all possible operators $M(t)$; rather, it formulates the minimal coarse-grained structure needed to derive self-similar spectral dynamics and scaling laws.
4.1 Log-shell partition and quadratic shell energies
Fix a ratio $q>1$ and define a logarithmic partition $\{\lambda_{\alpha}\}_{\alpha\in\mathbb{Z}}$ by
$\lambda_{\alpha}:=\lambda_{0}q^{\alpha},\qquad S_{\alpha}:=\{u:\ \lambda_{u}(t)\in[\lambda_{\alpha},\lambda_{\alpha+1})\},\qquad S_{\leq\alpha}:=\bigcup_{\gamma\leq\alpha}S_{\gamma}.$
We work with the quadratic shell energies introduced in Section 3.2:
$E_{\alpha}(t):=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}(t)^{2},\qquad E_{\leq\alpha}(t):=\sum_{\gamma\leq\alpha}E_{\gamma}(t).$
This choice is canonical because $\mathcal{L}(t)=\frac{1}{2}\|e_{t}\|_{2}^{2}=\sum_{\alpha}E_{\alpha}(t)$.
4.2 Exact shell bookkeeping and the role of dissipation
Section 3.4 already established the exact identity
$\frac{d}{dt}E_{\alpha}(t)=-D_{\alpha}(t)+\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t),\qquad D_{\alpha}(t):=\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2},$
where $\mathcal{F}_{\beta\to\alpha}(t)$ is the inter-shell quadratic-energy flux and satisfies $\mathcal{F}_{\beta\to\alpha}(t)=-\mathcal{F}_{\alpha\to\beta}(t)$. In particular, shell-internal coupling does not change $E_{\alpha}$ and the only mechanism that decreases $\sum_{\alpha}E_{\alpha}(t)$ is dissipation $\sum_{\alpha}D_{\alpha}(t)$.
4.3 Definition: renormalizable shell dynamics via cumulative energy flux
The notion of renormalizability adopted here follows the standard usage in statistical physics and field theory. Rather than requiring microscopic locality or exact continuum limits, external interaction effects on a resolution shell are integrated out and absorbed into effective inter-shell fluxes, while subleading corrections become irrelevant at coarse scales. This philosophy underlies Wilsonian renormalization group theory (wilson1983renormalization; kadanoff1966scaling), shell models of turbulence and energy cascades (kolmogorov1995turbulence), and effective hydrodynamic descriptions of nonequilibrium systems (forster1977large; spohn2012large). Our definition formalizes this principle at the level of spectral-shell energy dynamics.
Definition 4.1 (Renormalizable spectral-shell dynamics (energy form)).
The modewise dynamics are said to be (weakly) renormalizable with respect to the log-shell partition if, for sufficiently large $\alpha$, the shell energies admit a closed balance of the form
$\frac{d}{dt}E_{\alpha}(t)=-D_{\alpha}(t)+\mathcal{F}_{\alpha}^{\rm(net)}(t)+\mathcal{R}_{\alpha}(t),$
where the coarse-grained interaction contribution admits a net-flux form
$\mathcal{F}_{\alpha}^{\rm(net)}(t)=J_{\leq\alpha}(t)-J_{\leq\alpha-1}(t),$
and the remainder is negligible compared to the leading dissipation:
$|\mathcal{R}_{\alpha}(t)|\ll D_{\alpha}(t)\qquad\text{as }\alpha\to\infty.$
Here $J_{\leq\alpha}(t)$ is the cumulative quadratic-energy flux across the shell boundary at $\alpha$, defined by the exact microscopic inter-shell fluxes as
$\boxed{J_{\leq\alpha}(t):=\sum_{\beta>\alpha}\ \mathcal{F}_{\beta\to(\leq\alpha)}(t)=\sum_{\beta>\alpha}\ \sum_{\gamma\leq\alpha}\mathcal{F}_{\beta\to\gamma}(t).}$
Moreover, the cumulative flux is integrable at coarse scales in the sense that for each shell $\alpha$ there exists $T_{\alpha}>0$ such that for all $T>T_{\alpha}$,
$\int_{0}^{T}|J_{\leq\alpha}(t)|\,dt\;\leq\;C\int_{0}^{T}D_{\alpha}(t)\,dt.$
Remark 2 (What the cumulative flux represents).
Definition 4.1 does not assume microscopic locality of coupling. All shells $\beta>\alpha$ are included in $J_{\leq\alpha}$. Renormalizability asserts only that after coarse-graining the net inter-shell exchange can be summarized by a controlled boundary flux plus a negligible remainder.
Remark 3 (No direction is assumed).
Renormalizability imposes no restriction on the sign or direction of $J_{\leq\alpha}(t)$. All later results depend only on the existence of a controlled net-flux representation, not on an a priori cascade direction.
4.4 PDE-approximability across shells (energy continuum)
When the shell spacing is sufficiently fine over a resolution range, we may introduce an energy density $\varepsilon(\lambda,t)$ such that
$E_{\alpha}(t)\approx\int_{\lambda_{\alpha}}^{\lambda_{\alpha+1}}\varepsilon(\lambda,t)\,d\lambda,\qquad\mathcal{L}(t)=\int_{0}^{\infty}\varepsilon(\lambda,t)\,d\lambda.$
This is a local continuum approximation over that range; we do not require a global continuous-spectrum limit. As in nonequilibrium statistical physics forster1977large; spohn2012large, we do not assume the existence of an exact continuum limit; rather, renormalizability refers to the existence of a closed effective description after coarse-graining. We stress again that our continuum equations do not assume the existence of a genuine continuous-spectrum limit. They represent a local approximation of the discrete shell-flux difference when shells are sufficiently dense over a dynamically relevant resolution range. Formally, this amounts to approximating the discrete net-flux difference $J_{\leq\alpha}(t)-J_{\leq\alpha-1}(t)$ by a first-order finite-difference representation of $\partial_{\lambda}J$ over a dense but finite shell range.
4.5 Assumption: effective power-law transport (energy-flux interface)
Assumption 1 (Effective power-law transport (energy form)).
There exists a dynamically relevant resolution range $\lambda\in[\lambda_{\min},\lambda_{\max}]$ and a time-dependent amplitude $c(t)$ such that the renormalized energy flux can be expressed as
$J(\lambda,t)=v(\lambda,t)\,\varepsilon(\lambda,t),\qquad v(\lambda,t)\sim-c(t)\lambda^{b},\quad b>0,$
up to subleading corrections consistent with Definition 4.1. Equivalently, in the same range the coarse-grained energy dynamics admit an effective transport–dissipation form
$\partial_{t}\varepsilon(\lambda,t)+\partial_{\lambda}\!\big(v(\lambda,t)\varepsilon(\lambda,t)\big)=-2\lambda\,\varepsilon(\lambda,t),$
to leading order (the factor $2$ reflects quadratic energy dissipation).
Remark 4 (Scope).
Assumption 1 summarizes the outcome of coarse-graining rather than a microscopic property of $M(t)$. Providing operator-level conditions ensuring renormalizability and effective power-law transport is deferred to companion work.
5 Renormalized Energy-Shell Dynamics and the GRSD Regime
In Sections 34, we derived the exact modewise dynamics, introduced logarithmic spectral shells, and identified the quadratic shell energy $E_{\alpha}(t)=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}(t)^{2}$ as the canonical conserved quantity for shell-internal coupling. We now formulate the macroscopic regime of interest: a high-resolution tail where the inter-shell exchange is renormalizable and can be summarized by a controlled net-flux across shell boundaries.
5.1 The closed shell-level bookkeeping
Recall the exact shell balance law (Section 3.4):
$\frac{d}{dt}E_{\alpha}(t)=-D_{\alpha}(t)+\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t),\qquad D_{\alpha}(t):=\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2},$ (4)
where $\mathcal{F}_{\beta\to\alpha}(t)=-\mathcal{F}_{\alpha\to\beta}(t)$ is the inter-shell quadratic-energy exchange. In particular,
$\sum_{\alpha}\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t)=0,\qquad\frac{d}{dt}\Big(\sum_{\alpha}E_{\alpha}(t)\Big)=-\sum_{\alpha}D_{\alpha}(t)\leq 0.$
Thus, only dissipation decreases the total loss, while coupling redistributes energy across shells.
5.2 Renormalizable shell dynamics (energy form)
We work in the macroscopic closure regime postulated in Definition 4.1. Concretely, for sufficiently large $\alpha$ we assume the inter-shell exchange can be summarized as a controlled net-flux across the shell boundary $\alpha$ plus a negligible remainder:
$\frac{d}{dt}E_{\alpha}(t)=-D_{\alpha}(t)+\big(J_{\leq\alpha}(t)-J_{\leq\alpha-1}(t)\big)+\mathcal{R}_{\alpha}(t),\qquad|\mathcal{R}_{\alpha}(t)|\ll D_{\alpha}(t).$ (5)
This is a shell-level statement: it does not require a continuum limit, and it does not assume microscopic locality of coupling. It only asserts that after coarse-graining, the net exchange across shell boundaries is well-defined and controlled.
5.3 Continuum approximation (energy density)
Over a resolution range where shells are sufficiently fine, we may introduce an energy density $\varepsilon(\lambda,t)$ such that
$E_{\alpha}(t)\approx\int_{\lambda_{\alpha}}^{\lambda_{\alpha+1}}\varepsilon(\lambda,t)\,d\lambda,\qquad\mathcal{L}(t)=\sum_{\alpha}E_{\alpha}(t)\approx\int_{0}^{\infty}\varepsilon(\lambda,t)\,d\lambda.$
In this regime, the discrete net-flux difference $J_{\leq\alpha}-J_{\leq\alpha-1}$ may be approximated by a transport divergence $\partial_{\lambda}J(\lambda,t)$. Here “sufficiently fine” refers only to the scale separation between neighboring shells relative to the variation of the cumulative flux, not to any limiting process in which the shell spacing tends to zero.
5.4 Effective power-law transport (energy-flux interface)
We now specialize to the single effective interface used later (Assumption 1):
$J(\lambda,t)=v(\lambda,t)\,\varepsilon(\lambda,t),\qquad v(\lambda,t)\sim-c(t)\lambda^{b},\quad b>0,$ (6)
which yields, to leading order, the energy transport–dissipation equation
$\partial_{t}\varepsilon(\lambda,t)+\partial_{\lambda}\!\big(v(\lambda,t)\varepsilon(\lambda,t)\big)=-2\lambda\,\varepsilon(\lambda,t),$ (7)
on the relevant resolution range. The factor $2$ reflects the fact that dissipation in the mode ODE is $\lambda_{u}g_{u}$, hence quadratic energy dissipates as $\frac{d}{dt}\frac{1}{2}g_{u}^{2}=\lambda_{u}g_{u}^{2}$.
5.5 RMS amplitude notation
For compatibility with earlier GSD/GRSD forms, we also define an RMS amplitude density
$g(\lambda,t):=\sqrt{2\,\varepsilon(\lambda,t)}.$ (8)
Then
$\mathcal{L}(t)=\int_{0}^{\infty}\varepsilon(\lambda,t)\,d\lambda=\frac{1}{2}\int_{0}^{\infty}g(\lambda,t)^{2}\,d\lambda,$
and equation (7) can be viewed as the energy-level backbone behind the GRSD tail forms written in terms of $g$.
5.6 Summary
The key structural takeaway is:
Shell-internal coupling is a rotation that preserves quadratic energy within each shell. Any macroscopic evolution of shell energies is fully captured by (i) dissipation and (ii) inter-shell boundary flux.
All scaling-law consequences in later sections are derived from the renormalized boundary-flux interface (5)–(6) in a high-resolution range.
6 Self-Similar GRSD Tail Under Power-Law Drift
In this section we analyze the effective GRSD regime implied by the energy transport interface (7) with power-law drift $v(\lambda,t)\sim-c(t)\lambda^{b}$. We emphasize that the discrete shell bookkeeping is primary; the continuum calculation below is an approximation valid over a dense shell range. Equivalently, all results in this section can be interpreted as asymptotic solutions of the discrete shell-energy dynamics, with the PDE serving only as a convenient analytic representation.
6.1 Energy transport equation and characteristics
We start from the leading-order energy PDE
$\partial_{t}\varepsilon+\partial_{\lambda}(v\varepsilon)=-2\lambda\,\varepsilon,\qquad v(\lambda,t)\sim-c(t)\lambda^{b}.$
Expanding the transport term gives
$\partial_{t}\varepsilon+v\,\partial_{\lambda}\varepsilon+(\partial_{\lambda}v)\,\varepsilon=-2\lambda\,\varepsilon.$ (9)
Along characteristics $\lambda(t)$ satisfying
$\frac{d\lambda}{dt}=v(\lambda,t)\sim-c(t)\lambda^{b},$ (10)
the energy density evolves by
$\frac{d}{dt}\log\varepsilon(\lambda(t),t)=-\partial_{\lambda}v(\lambda(t),t)-2\lambda(t).$ (11)
Define the effective training time
$\tau(t):=\int_{0}^{t}c(s)\,ds.$ (12)
For $b\neq 1$, the characteristic flow integrates to
$\lambda(t)^{1-b}=\lambda_{0}^{1-b}+(b-1)\tau(t),\qquad(b\neq 1).$ (13)
In the supercritical regime $b>1$, characteristics move from high to lower resolution progressively, enabling a resolution cascade over time.
6.2 Universal tail factor from transport Jacobian
Using $\partial_{\lambda}v(\lambda,t)\sim-bc(t)\lambda^{b-1}$, equation (11) becomes
$\frac{d}{dt}\log\varepsilon(\lambda(t),t)=bc(t)\lambda(t)^{b-1}-2\lambda(t).$
The first term is purely geometric: it comes from the Jacobian of the spectral flow (compression/expansion in $\lambda$), while the second term is physical dissipation.
Ignoring dissipation temporarily, the transport-only equation $\partial_{t}\varepsilon+\partial_{\lambda}(v\varepsilon)=0$ yields the universal scaling
$\varepsilon(\lambda,t)\propto\lambda^{-b}.$ (14)
Equivalently, the RMS amplitude density satisfies the transport-only tail
$g(\lambda,t)\propto\lambda^{-b/2}.$ (15)
This tail exponent is fixed by the power-law drift and does not depend on the microscopic details of $M(t)$, only on the renormalized transport interface.
6.3 Dissipation and the resolution frontier
We now account for dissipation. Along a characteristic, define the dissipation action
$\Phi(\lambda,t):=\int_{0}^{t}2\lambda(s)\,ds,$ (16)
where $\lambda(s)=\lambda(s;\lambda_{0})$ follows (10). In the supercritical regime $b>1$, inserting the asymptotic behavior $\lambda(t)\sim\big((b-1)\tau(t)\big)^{-1/(b-1)}$ into (16) yields the scaling
$\Phi(\lambda,t)\;\asymp\;K\,\lambda^{\,b-1}\,\tau(t),$ (17)
for a constant $K>0$ (the exact prefactor is not important for scaling).
This produces a sharp cutoff in the energy density: modes with $\lambda^{b-1}\tau(t)\gg 1$ are exponentially suppressed. Hence the dynamics localize around a resolution frontier
$\lambda^{\ast}(t)\;\asymp\;\tau(t)^{-1/(b-1)}.$ (18)
6.4 GRSD tail form (energy and amplitude)
Combining the universal transport tail (14) with the dissipation action (17), we obtain the GRSD tail template in energy form:
$\boxed{\varepsilon(\lambda,t)=A_{\varepsilon}(t)\,\lambda^{-b}\,\exp\!\big(-K\,\lambda^{\,b-1}\tau(t)\big),}$ (19)
where $A_{\varepsilon}(t)$ is a slowly varying (typically polynomial) amplitude determined by matching to lower-resolution shells / initial data.
Equivalently, in RMS amplitude notation $g=\sqrt{2\varepsilon}$,
$\boxed{g(\lambda,t)=A_{g}(t)\,\lambda^{-b/2}\,\exp\!\Big(-\tfrac{1}{2}K\,\lambda^{\,b-1}\tau(t)\Big),\qquad A_{g}(t):=\sqrt{2A_{\varepsilon}(t)}.}$ (20)
6.5 Remarks on regimes $b\leq 1$
The critical case $b=1$ corresponds to exponential-in-$\tau$ characteristics and does not produce a progressive resolution frontier. The subcritical case $0<b<1$ reaches the resolution floor in finite $\tau$ and therefore cannot sustain asymptotic data scaling. Since modern scalable training empirically exhibits progressive refinement, we focus on $b>1$ in the main text and discuss $b\leq 1$ in Appendix LABEL:app:b<1.
6.6 Summary
Under the renormalized shell-flux interface and effective power-law drift, the high-resolution tail admits a universal GRSD form: a power-law transport tail multiplied by an exponential dissipation cutoff, with a resolution frontier $\lambda^{\ast}(t)\asymp\tau(t)^{-1/(b-1)}$.
7 Scaling Laws and the Geometry of Double Descent
We now derive scaling-law behavior from the GRSD template. The correct global loss is the total quadratic error energy
$\mathcal{L}(t)=\frac{1}{2}\|e_{t}\|_{2}^{2}=\sum_{\alpha}E_{\alpha}(t)\approx\int_{0}^{\infty}\varepsilon(\lambda,t)\,d\lambda=\frac{1}{2}\int_{0}^{\infty}g(\lambda,t)^{2}\,d\lambda.$
Thus the scaling analysis should be performed on $\varepsilon(\lambda,t)$ (or equivalently $g^{2}$).
7.1 Loss scaling from the GRSD tail
Insert the energy-form GRSD template (19):
$\varepsilon(\lambda,t)=A_{\varepsilon}(t)\lambda^{-b}\exp\!\big(-K\lambda^{b-1}\tau(t)\big).$
For large $\tau(t)$, the integral is dominated near the frontier $\lambda^{\ast}(t)\asymp\tau(t)^{-1/(b-1)}$. A standard Laplace-type estimate yields
$\mathcal{L}(t)\;\asymp\;A_{\varepsilon}(t)\,\big(\lambda^{\ast}(t)\big)^{1-b}.$ (21)
Using $\lambda^{\ast}(t)\asymp\tau(t)^{-1/(b-1)}$ gives
$\mathcal{L}(t)\;\asymp\;A_{\varepsilon}(t)\,\tau(t)^{\frac{b-1}{b-1}\cdot\frac{b-1-(1)}{\,b-1\,}}=A_{\varepsilon}(t)\,\tau(t)^{\frac{b-1-(1)}{\,b-1\,}}=A_{\varepsilon}(t)\,\tau(t)^{-\frac{b-2}{\,b-1\,}}.$ (22)
Hence, once $A_{\varepsilon}(t)$ is controlled (typically polynomial/slowly varying), the dominant exponent is determined by $b$.
Remark 5 (Where $A_{\varepsilon}(t)$ comes from).
$A_{\varepsilon}(t)$ is not an arbitrary free function: it is fixed by matching the high-resolution tail to the lower-resolution part of the spectrum and by conservation/dissipation constraints at the shell level. In particular, $A_{\varepsilon}(t)$ encodes how much energy is injected into the tail by inter-shell transport at a given effective time. Our framework isolates the universal tail shape; the prefactor can be estimated empirically from spectral measurements.
7.2 Double Descent from In-Domain Finite-Sample Spectral Mismatch
We now explain the origin of the double-descent phenomenon from the perspective of renormalizable spectral-shell dynamics. Crucially, the mechanism described here is entirely in-domain: training and test data are drawn independently from the same underlying data-generating distribution. The non-monotonicity of the test loss arises not from distribution shift, but from unavoidable finite-sample spectral mismatch concentrated in high-resolution modes.
Training dynamics and monotone training loss.
Throughout training, the error evolves according to
$\dot{e}_{t}=-M(t)e_{t},\qquad M(t)=J_{\theta(t)}J_{\theta(t)}^{\!*}\succeq 0.$
Measured in the training metric, the quadratic loss
$\mathcal{L}_{\rm tr}(t):=\frac{1}{2}\|e_{t}\|_{L^{2}(p)}^{2}=\sum_{\alpha}E_{\alpha}(t),\qquad E_{\alpha}(t):=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}(t)^{2},$
is strictly non-increasing:
$\frac{d}{dt}\mathcal{L}_{\rm tr}(t)=-\langle e_{t},M(t)e_{t}\rangle\leq 0.$
This monotonicity is exact and does not rely on any coarse-graining, continuum approximation, or weak-coupling assumption.
In-domain train/test mismatch as a finite-sample effect.
Let $P_{\rm tr}$ and $P_{\rm te}$ denote the empirical quadratic forms induced by the training and test samples, respectively. Both are unbiased estimators of the same population operator associated with the underlying distribution, but differ due to finite sampling. Define the mismatch operator
$\Delta:=P_{\rm te}-P_{\rm tr}.$
The test loss can be written as
$\mathcal{L}_{\rm te}(t)=\frac{1}{2}\langle e_{t},P_{\rm te}e_{t}\rangle=\mathcal{L}_{\rm tr}(t)+\frac{1}{2}\langle e_{t},\Delta e_{t}\rangle.$
The second term captures purely finite-sample effects and vanishes only in the infinite-data limit.
For continuous data distributions or infinite-rank kernel operators, the population spectrum exhibits vanishing eigenvalue gaps in the spectral tail. Classical results on empirical covariance and integral operators show that, in this regime, the associated empirical eigenspaces exhibit $O(1)$ fluctuations under finite sampling (koltchinskii2017normal; rosasco2010learning). Consequently, even under in-domain sampling, $\Delta$ is generically supported on high-resolution (spectral-tail) components and is not co-diagonalizable with the training-induced operator $M(t)$.
Shell amplification via geometric growth of $A(t)$.
The effect of this mismatch becomes dynamically visible through the shell structure of the error. Under the renormalizable shell-dynamics assumption, the high-resolution tail of the error admits the generalized resolution spectral form
$g(\lambda,t)\;\approx\;A(t)\,\lambda^{-b}\,e^{-\phi(\lambda,t)},\qquad b>1,$
over a dynamically relevant range of resolutions. As shown in Section LABEL:sec:weak_coupling, the geometric prefactor $A(t)$ grows polynomially during an initial training phase, reflecting compression of spectral coordinates induced by transport across resolution shells.
Since the mismatch operator $\Delta$ is concentrated in high-resolution shells, the correction term satisfies a lower bound of the form
$\langle e_{t},\Delta e_{t}\rangle\;\gtrsim\;\sum_{\alpha\geq\alpha_{0}}\delta_{\alpha}\,E_{\alpha}(t),\qquad\delta_{\alpha}>0,$
for some tail index $\alpha_{0}$. In the early-time regime where dissipation is not yet active on the tail, the shell energies scale as
$\sum_{\alpha\geq\alpha_{0}}E_{\alpha}(t)\;\asymp\;A(t)^{2}\int_{\lambda_{\alpha_{0}}}^{\infty}\lambda^{-2b}\,d\lambda,$
so that the test correction grows proportionally to $A(t)^{2}$. As a result, the growth of $A(t)$ can temporarily dominate the monotone decay of $\mathcal{L}_{\rm tr}(t)$, forcing the test loss $\mathcal{L}_{\rm te}(t)$ to increase even though training loss continues to decrease.
Dissipation, resolution frontier, and eventual decay.
At later times, dissipation encoded by the factor $e^{-\phi(\lambda,t)}$ suppresses the spectral tail and localizes the error near the resolution frontier $\lambda^{*}(t)$. Both training and test losses then decay to zero, and the finite-sample mismatch term becomes negligible. The test loss therefore exhibits a rise–then–fall profile: an initial increase driven by shell-level geometric amplification, followed by universal decay once dissipation dominates.
Interpretation.
Double descent thus emerges as a generic consequence of three ingredients: (i) exact monotonic dissipation in the training metric, (ii) unavoidable finite-sample spectral mismatch in high-resolution modes, and (iii) geometric amplification of tail energy induced by spectral-shell transport. This mechanism operates entirely in-domain and does not rely on distribution shift, overparameterization heuristics, or ad hoc regularization. Instead, it reflects a structural interaction between sampling noise and the renormalizable shell dynamics governing neural training.
7.3 Unified view of NTK and feature learning through spectral transport
The shell-energy formulation provides a clean unification of lazy (NTK-like) training and feature learning through a single macroscopic quantity: the spectral drift velocity $v(\lambda,t)$. Recall the definition
$J(\lambda,t)=v(\lambda,t)\,g(\lambda,t),$
which defines the effective transport velocity $v(\lambda,t)$ in the coarse-grained description.
Lazy / NTK regime ($v\equiv 0$).
If the kernel operator is effectively frozen during training ($\dot{M}(t)\approx 0$), then $v(\lambda,t)\equiv 0$ and no inter-shell transport occurs. The shell dynamics reduce to
$\frac{d}{dt}E_{\alpha}(t)=-D_{\alpha}(t),$
so each shell decays independently under dissipation. No resolution frontier forms, and the dynamics coincide with classical NTK theory.
Feature learning regime ($v\neq 0$).
When the representation evolves ($\dot{M}(t)\neq 0$), the drift velocity becomes active and induces inter-shell energy transport. Energy is redistributed across resolutions before being dissipated, producing a moving resolution frontier and the GRSD tail described in Sections 5–6.
Crucially, this redistribution preserves total quadratic energy and does not increase training loss. Its observable consequences depend on how the transported energy aligns with the evaluation distribution (training vs. test).
Continuum of regimes.
Between these extremes lies a continuum:
  • If $|v(\lambda,t)|\ll\lambda$, dissipation dominates and training is effectively NTK-like.
  • If $|v(\lambda,t)|$ is comparable to or larger than $\lambda$ over a resolution range, transport reshapes the spectrum and feature learning emerges.
Thus, lazy training and feature learning are not distinct dynamical theories, but limiting behaviors of the same renormalized shell-energy dynamics.
Interpretation.
From this perspective, feature learning corresponds to spectral energy transport induced by representation drift, while NTK corresponds to the degenerate zero-transport limit. Both are unified within the same operator-theoretic framework.
7.4 Summary
In the GRSD regime ($b>1$), the renormalized energy-shell dynamics yield: (i) a universal high-resolution tail shape (19), (ii) a resolution frontier $\lambda^{\ast}(t)\asymp\tau(t)^{-1/(b-1)}$, and (iii) scaling-law decay of the total loss controlled by the drift exponent $b$ and the effective time $\tau(t)$.
8 Related Work
Neural scaling laws.
Empirical studies have established that neural networks display remarkably regular power-law relationships between compute, model size, dataset size, and achievable loss (kaplan2020scaling; henighan2020scaling; hestness2017deep; hoffmann2022training; hernandez2021scaling). Recent work has investigated both the emergence of such scaling behavior and its theoretical underpinnings, including dynamical models of loss evolution (bordelon2024dynamical) and the role of data pruning (sorscher2022beyond). Precision scaling laws have also been explored in the context of architectural and quantization constraints (kumar2024scaling).
Feature learning and spectral bias.
A major line of work studies how neural networks acquire hierarchical representations and exhibit spectral preference (rahaman2019spectral; bordelon2023feature). Kernel-based analyses of learning dynamics (bietti2021inductive; canatar2022kernel) and infinite-width approximations (jacot2018neural; lee2019wide; yang2021tensor) have contributed significantly to understanding the transition between lazy training and representation learning. Recent studies have also revealed consistency of learned features across widths (vyas2023feature) and the spectral evolution of networks (wang2023spectralevolution).
Optimization dynamics and stability.
The geometry of optimization landscapes, the effect of batch size, and training stability have been examined extensively (keskar2017large; ghorbani2019investigation). Work on pruning and compression (rosenfeld2021predictability; han2015deep; blalock2020state; lecun1990optimal; molchanov2017variational; lee2019snip; wang2020picking; han2016deep; nagel2021up; frantar2022gptq; dettmers2022llm) has illuminated how spectral structure interacts with parameter sparsity and low-precision computation. Theoretical connections between spectral evolution, implicit bias, and neural dynamics continue to be an active area of research (domine2024lazy; zhang2025generalized).
Perturbation theory and adiabatic analysis.
Our spectral formulation draws on classical operator perturbation theory (kato2012short) and its analogues in quantum adiabatic evolution (Zwiebach_Adiabatic). These tools formalize how eigenfunctions and eigenvalues evolve under smooth or weakly coupled updates, providing a principled foundation for the drift–dissipation dynamics developed in this work.
9 Conclusion
This work proposed a spectral–shell framework for understanding neural scaling laws, feature learning, and double-descent phenomena directly from the operator-level dynamics induced by gradient descent. Starting from the exact function-space evolution $\dot{e}_{t}=-M(t)e_{t}$, we derived a rigorous modewise formulation and showed that, upon logarithmic coarse-graining, shell-internal interactions cancel identically at the level of quadratic error energy. As a result, the macroscopic evolution of error is governed entirely by dissipation and inter-shell energy exchange.
The central modeling ingredient of this paper is a renormalizable spectral-shell dynamics assumption: after coarse-graining, the cumulative effect of microscopic interactions can be summarized by a controlled net energy flux across shell boundaries. Under an effective power-law form of this renormalized shell-level flux, the shell dynamics admit a self-similar high-resolution tail with a moving resolution frontier. This structure yields explicit scaling-law behavior for the total loss and provides a unified explanation of neural scaling laws and double descent.
A key conceptual outcome of this framework is a unified view of lazy (NTK-like) training and feature learning. When inter-shell transport is negligible, shell energies decay independently under dissipation, recovering classical kernel dynamics. When transport is active, spectral energy is redistributed across resolutions before being dissipated, inducing representation learning. Both regimes—and the continuum between them—are governed by the same shell-energy bookkeeping, differing only in the effective shell-flux interface.
Importantly, the analysis does not assume the existence of a continuous spectrum or a vanishing shell spacing. All continuum partial differential equations appearing in the paper serve only as convenient local approximations of discrete shell-energy dynamics over resolution ranges where shells are sufficiently dense. All scaling-law conclusions can be equivalently interpreted at the discrete shell level.
Future directions.
Several important questions remain open:
  1. 1.
    Multi-task and multi-distribution learning. The present analysis focuses on a single task and a single data distribution. Extending spectral-shell dynamics to multi-task settings—where different task operators may not share eigenbases and may interact through shared representation drift—could reveal new mechanisms of transfer, interference, and modular generalization.
  2. 2.
    Origin of renormalizability and power-law shell transport. In this work, the renormalizable shell-flux interface and its effective power-law form are treated as macroscopic assumptions. A natural next step is to understand their qualitative and quantitative origins from operator-level structure. Possible mechanisms include locality of spectral coupling, effective one-directional energy transfer across resolution shells, and the suppression of long-range interactions by stochastic optimization noise. Clarifying when and why such mechanisms produce renormalizable and power-law shell dynamics would substantially deepen the theoretical foundations of the framework.
  3. 3.
    Modeling optimizers and learning-rate schedules. The present formulation is developed for gradient-based optimization with an effective time reparameterization. Momentum, adaptive methods, and second-order or approximate natural-gradient schemes can modify both dissipation and transport. Developing spectral-shell descriptions of these optimizers—and of common learning-rate schedules—may explain optimizer-dependent variations in scaling behavior.
  4. 4.
    Connecting $J_{\theta}$ to network architecture. The Jacobian operator $J_{\theta}$ mediates how parameters generate spectral structure in function space. Understanding how architectural features such as depth, width, and parameterization shape the induced shell dynamics could connect model scaling, representational capacity, and achievable loss within a unified operator-theoretic framework.
Overall, the spectral-shell perspective developed here isolates a minimal and robust macroscopic structure underlying neural training dynamics. By separating exact shell-level conservation laws from coarse-grained flux assumptions, it provides a flexible foundation for understanding scaling phenomena while leaving room for future work on their microscopic origins.
Appendix A Exact Shell-Energy Bookkeeping from Modewise Dynamics
This appendix derives the exact shell-level energy balance underlying the main text. Its purpose is purely structural: to show how the quadratic shell energies used throughout the paper arise rigorously from the exact modewise dynamics. No weak-coupling, locality, or continuum assumptions are imposed here. All scaling-law results rely only on the renormalizable shell-dynamics assumption stated in Definition 4.1.
A.1 Modewise dynamics and antisymmetric coupling
Recall from Section 3.1 that, in the instantaneous eigenbasis of $M(t)$, the error amplitudes satisfy the exact system
$\partial_{t}g_{u}(t)+\sum_{v\neq u}g_{v}(t)\,\Omega_{v\to u}(t)=-\lambda_{u}(t)\,g_{u}(t),$ (23)
with coupling coefficients
$\Omega_{v\to u}(t)=\frac{\langle\varphi_{u}(t),\dot{M}(t)\varphi_{v}(t)\rangle}{\lambda_{v}(t)-\lambda_{u}(t)},\qquad u\neq v.$
Since $\dot{M}(t)$ is self-adjoint, these coefficients satisfy the strict antisymmetry
$\Omega_{v\to u}(t)=-\Omega_{u\to v}(t),\qquad u\neq v.$ (24)
A.2 Logarithmic shells and quadratic energies
Fix a logarithmic partition of the spectrum as in Section 3.2:
$S_{\alpha}:=\{u:\ \lambda_{u}(t)\in[\lambda_{\alpha},\lambda_{\alpha+1})\},\qquad\lambda_{\alpha}=\lambda_{0}q^{\alpha},\quad q>1.$
For each shell, define the quadratic shell energy
$E_{\alpha}(t):=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}(t)^{2}.$ (25)
This choice is canonical, since the total training loss is exactly
$\mathcal{L}(t)=\frac{1}{2}\|e_{t}\|_{2}^{2}=\sum_{\alpha}E_{\alpha}(t).$
A.3 Exact cancellation of shell-internal coupling
Differentiate (25):
$\frac{d}{dt}E_{\alpha}(t)=\sum_{u\in S_{\alpha}}g_{u}(t)\,\partial_{t}g_{u}(t).$
Substituting (23) yields
$\frac{d}{dt}E_{\alpha}(t)=-\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2}-\sum_{u\in S_{\alpha}}\sum_{v\neq u}g_{u}(t)g_{v}(t)\,\Omega_{v\to u}(t).$ (26)
Split the coupling sum into contributions with $v\in S_{\alpha}$ and $v\notin S_{\alpha}$. By antisymmetry (24), the shell-internal contribution vanishes identically:
$\sum_{u\in S_{\alpha}}\sum_{\begin{subarray}{c}v\in S_{\alpha}\\ v\neq u\end{subarray}}g_{u}g_{v}\,\Omega_{v\to u}=0.$ (27)
Thus shell-internal coupling acts as a pure rotation of amplitudes and does not change the quadratic shell energy.
A.4 Exact shell-energy balance
Only dissipation and inter-shell exchange remain. Define the inter-shell quadratic-energy flux
$\mathcal{F}_{\beta\to\alpha}(t):=-\sum_{\begin{subarray}{c}u\in S_{\alpha}\\ v\in S_{\beta}\end{subarray}}g_{u}(t)g_{v}(t)\,\Omega_{v\to u}(t).$
Then (26) reduces to the exact identity
$\boxed{\frac{d}{dt}E_{\alpha}(t)=-\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2}+\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t).}$ (28)
Moreover, antisymmetry implies the action–reaction property
$\mathcal{F}_{\beta\to\alpha}(t)=-\mathcal{F}_{\alpha\to\beta}(t),$
so that inter-shell coupling conserves total quadratic energy:
$\sum_{\alpha}\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t)=0.$
The only mechanism that decreases $\sum_{\alpha}E_{\alpha}(t)$ is dissipation.
A.5 Relation to renormalizable shell dynamics
Equation (28) is an exact bookkeeping identity. It involves no approximation, no continuum limit, and no assumption on the structure of the coupling coefficients beyond antisymmetry.
The renormalizable shell-dynamics assumption introduced in Definition 4.1 is a macroscopic closure of this identity: it postulates that, after coarse-graining, the cumulative effect of the inter-shell fluxes can be summarized by a controlled net boundary flux plus a negligible remainder. All scaling-law results in the main text are derived from this shell-level closure and do not rely on any further microscopic characterization.