### v4

# Renormalizable Spectral-Shell Dynamics as the Origin of Neural Scaling Laws

Abstract
Neural scaling laws and double-descent phenomena suggest that deep-network training obeys a simple macroscopic structure despite highly nonlinear optimization dynamics. We derive such structure directly from gradient descent in function space. For mean-squared error loss, the training error evolves as $\dot{e}_{t}=-M(t)e_{t}$ with $M(t)=J_{\theta(t)}J_{\theta(t)}^{\!*}$, a time-dependent self-adjoint operator induced by the network Jacobian. Using Kato perturbation theory, we obtain an exact system of coupled modewise ODEs in the instantaneous eigenbasis of $M(t)$.
To extract macroscopic behavior, we introduce a logarithmic spectral-shell coarse-graining and track quadratic error energy across shells. Microscopic interactions within each shell cancel identically at the energy level, so shell energies evolve only through dissipation and external inter-shell interactions. We formalize this via a renormalizable shell-dynamics assumption, under which cumulative microscopic effects reduce to a controlled net flux across shell boundaries. This shell-form equation unifies lazy-regime training and feature learning as two limits of the same spectral-shell dynamics.
Assuming an effective power-law spectral transport in a relevant resolution range, the shell dynamics admits a self-similar solution with a moving resolution frontier and explicit scaling exponents, which explains neural scaling laws and double descent.
1 Introduction
Modern deep neural networks are trained by strongly nonlinear and high-dimensional optimization procedures. Nevertheless, across architectures, datasets, and training recipes, their performance exhibits remarkably simple and robust regularities. Most prominently, empirical studies have shown that training and test losses obey approximate power laws in model size, dataset size, or total compute over wide dynamic ranges (hestness2017deep; kaplan2020scaling; henighan2020scaling; hoffmann2022training; hernandez2021scaling; wei2022emergent). Related law-like behaviors have been observed in precision scaling, pruning, sparsification, and model densification (kumar2024scaling; sorscher2022beyond; rosenfeld2021predictability; blalock2020state; Xiao2024DensingLaw). These phenomena strongly suggest the existence of an underlying macroscopic structure governing how error is redistributed and dissipated during training.
Limitations of existing theories.
A classical theoretical explanation is provided by the neural tangent kernel (NTK) framework (jacot2018ntk; lee2019wide; bietti2021inductive), which models training as a linear dynamics in function space with a fixed kernel. While this perspective explains certain spectral biases and convergence rates, it relies on a frozen feature map and breaks down once representation learning becomes significant. Recent work has therefore emphasized feature-learning dynamics, using mean-field, tensor-program, or dynamical systems approaches (yang2021tensor; canatar2022kernel; bordelon2024dynamical; bordelon2023feature; wang2023spectralevolution). These models successfully capture richer behavior beyond the lazy regime, but typically treat NTK-like and feature-learning regimes via distinct approximations and do not provide a single operator-level equation that interpolates continuously between them.
A spectral-shell perspective.
In this work, we take a function-space and operator-theoretic viewpoint. For mean-squared error loss, the training error $e_{t}$ evolves according to the exact linear equation
$\dot{e}_{t}=-M(t)e_{t},\qquad M(t)=J_{\theta(t)}J_{\theta(t)}^{\!*},$
where $M(t)$ is a time-dependent, self-adjoint operator induced by the network Jacobian. Rather than assuming fixed features or a prescribed kernel, we analyze the training dynamics directly through the evolving spectral structure of $M(t)$.
Using Kato perturbation theory, we derive an exact system of coupled modewise ODEs in the instantaneous eigenbasis of $M(t)$. The resulting dynamics are fully rigorous but highly nonlocal across modes. To extract macroscopic behavior, we introduce a logarithmic spectral-shell coarse-graining and track the evolution of quadratic error energy within each shell. A key structural fact is that microscopic interactions within a shell, arising from eigenbasis drift, are strictly antisymmetric and cancel identically at the level of quadratic energy. Consequently, shell energies evolve only through dissipation and external interaction effects—namely, energy exchange with other shells.
Renormalizable shell dynamics.
This observation leads to an exact shell-level energy balance law. To close it at macroscopic scales, we introduce a renormalizable shell-dynamics assumption: after coarse-graining, the cumulative effect of microscopic degrees of freedom can be absorbed into a controlled net flux across shell boundaries, with subleading corrections becoming negligible. This notion of renormalizability is inspired by analogous constructions in statistical physics and condensed-matter theory, and it does not require a continuum limit nor impose any a priori direction of spectral transport.
When spectral shells are sufficiently dense over a dynamically relevant range, the shell dynamics admits an continuum approximation in the spectral variable $\lambda$, yielding an effective transport–dissipation equation. We emphasize that this PDE is a convenient representation of shell dynamics, not its fundamental starting point. Throughout the paper, renormalizability and effective power-law transport are treated as macroscopic assumptions, not as consequences of weak coupling or locality.
Effective transport and scaling laws.
We further assume that, in the relevant resolution range, the renormalized spectral flux can be summarized by an effective power-law transport form. Under this single coarse-grained assumption, the shell dynamics admits a self-similar solution with a moving resolution frontier, geometric amplitude growth, and power-law dissipation. This structure yields explicit scaling-law exponents and provides a unified explanation for neural scaling laws and double-descent phenomena.
Within this framework, lazy (NTK-like) training and feature learning arise as two limits of the same spectral-shell dynamics. When the effective transport vanishes, the dynamics reduce to pure dissipation with fixed features. When transport is active, spectral mass flows across resolutions, inducing representation learning. Thus, both regimes—and the continuum between them—are unified by a single operator-level shell dynamics governing the redistribution of error during training.
2 Preliminaries: Function-Space Dynamics and Error Evolution
We consider supervised learning with mean-squared error (MSE) loss in the Hilbert space $H=L^{2}(p)$ induced by the data distribution $p(x)$. All functions are identified up to sets of measure zero, and the inner product is defined by
$\langle f,g\rangle:=\mathbb{E}_{x\sim p}\,[\,f(x)\,g(x)\,],\qquad\|f\|_{2}^{2}=\langle f,f\rangle.$
2.1 Neural networks as functions in $L^{2}(p)$
A neural network with parameters $\theta\in\mathbb{R}^{N}$ represents a time-dependent function $f_{\theta(t)}\in H$. Given a target function $f^{*}\in H$, we define the error at time $t$ as
$e_{t}:=f^{*}-f_{\theta(t)}.$
Throughout this work, we emphasize that no structural assumptions such as fixed features, invariant kernels, or predetermined mode decompositions are imposed: the representation $f_{\theta(t)}$ is allowed to evolve freely in $H$, and all structure in the dynamics arises directly from gradient descent.
2.2 Gradient descent induces a linear evolution in function space
For MSE loss,
$\mathcal{L}(\theta)=\frac{1}{2}\,\|f^{*}-f_{\theta}\|_{2}^{2}=\frac{1}{2}\,\|e\|_{2}^{2},$
the gradient flow in parameter space is
$\dot{\theta}(t)=-\nabla_{\theta}\mathcal{L}(\theta(t)).$
By the chain rule,
$\dot{e}_{t}=-J_{\theta(t)}\,\dot{\theta}(t),$
where $J_{\theta}\colon\mathbb{R}^{N}\to H$ denotes the Jacobian operator $(J_{\theta}v)(x)=\nabla_{\theta}f_{\theta}(x)\cdot v$. Substituting the gradient flow equation gives
$\dot{e}_{t}=-J_{\theta(t)}\,J_{\theta(t)}^{\!*}\,e_{t}.$
Thus the error evolves according to a linear operator in function space:
$\boxed{\dot{e}_{t}=-M(t)\,e_{t},\qquad M(t):=J_{\theta(t)}\,J_{\theta(t)}^{\!*}.}$
The operator $M(t)$ is self-adjoint and positive semidefinite. It is the only operator that governs the evolution of the error in our analysis; its time dependence reflects the evolving representation of the network under gradient descent.
In this section we work in the continuous-time gradient flow limit,
$\dot{\theta}(t)=-\nabla_{\theta}\mathcal{L}(\theta(t)),$
which corresponds to taking the learning rate to be infinitesimal. A discrete-time update with (possibly time-varying) step sizes $\eta_{k}$,
$\theta_{k+1}=\theta_{k}-\eta_{k}\nabla_{\theta}\mathcal{L}(\theta_{k}),$
can be viewed as a time reparameterization of this flow, where the effective training time is proportional to the accumulated step size $\sum_{k}\eta_{k}$. Thus, learning rate schedules do not change the form of the operator dynamics $\dot{e}_{t}=-M(t)e_{t}$; instead, they induce a non-uniform rescaling of time that will later be absorbed into the effective time variable $\tau(t)$ in our spectral analysis.
2.3 Spectral decomposition of $M(t)$
For each fixed $t$, the operator $M(t)$ acts on a finite-dimensional subspace of $H$ determined by the network’s Jacobian, and therefore admits a discrete spectral decomposition. We write
$M(t)\,\varphi_{u}(t)=\lambda_{u}(t)\,\varphi_{u}(t),\qquad u\in\mathcal{U}(t),$
where $\mathcal{U}(t)$ is a finite or countable index set, the eigenvalues $\lambda_{u}(t)\geq 0$, and the eigenfunctions $\{\varphi_{u}(t)\}_{u\in\mathcal{U}(t)}$ form an orthonormal family in $H$. Expanding the error in this moving eigenbasis gives
$e_{t}=\sum_{u\in\mathcal{U}(t)}g_{u}(t)\,\varphi_{u}(t).$
The amplitudes $g_{u}(t)$ encode how much error lies along each instantaneous mode of $M(t)$, and the time dependence of both $\lambda_{u}(t)$ and $\varphi_{u}(t)$ reflects the evolution of the network’s representation.
2.4 Why a spectral formulation?
Although $M(t)$ generally has low rank compared to the ambient dimension of $H$, its eigenstructure provides a natural lens on the error dynamics. The eigenvalues $\lambda_{u}(t)$ quantify the rate at which error aligned with mode $u$ is dissipated, while the evolution of $\varphi_{u}(t)$ captures the “feature learning” aspect of training. In Section 3.1, we show that the amplitudes $g_{u}(t)$ satisfy a rigorous, coupled system of mode ODEs derived from Kato’s perturbation theory. In Section 3, we then coarse-grain these discrete modes to obtain a continuous spectral PDE that describes the statistical redistribution of error across many nearby modes.
3 From Modewise ODEs to Coarse-Grained Spectral Dynamics
In this section we derive the fundamental coarse-grained spectral structure governing the dynamics of the error under gradient descent. Starting from the exact function-space evolution $\dot{e}_{t}=-M(t)e_{t}$, we first express the dynamics in the instantaneous eigenbasis of $M(t)$, obtaining a fully rigorous system of coupled modewise ODEs.
Rather than attempting to track individual modes, which is intractable at macroscopic scales, we then introduce a logarithmic spectral-shell coarse-graining and study the evolution of the quadratic error energy carried by each shell. This leads to an exact shell-level balance law, in which shell-internal coupling cancels identically and all nontrivial interactions appear as inter-shell energy fluxes.
When the shell spacing is sufficiently fine over a dynamically relevant range, this shell dynamics admits an continuum approximation in the spectral variable $\lambda$, yielding a transport–dissipation PDE. However, the shell-level formulation itself is primary and does not rely on any continuum limit. In particular, we never assume that the spectrum becomes continuous or that a limit as the shell spacing vanishes exists; all continuum expressions should be read as local approximations to discrete shell differences.
3.1 Exact Mode ODEs in a Drifting Eigenbasis
For each time $t$, the operator $M(t)=J_{\theta(t)}J_{\theta(t)}^{\!*}$ is self-adjoint and positive semidefinite, acting on a finite-dimensional subspace of $H=L^{2}(p)$. Thus its spectrum is discrete. Let
$M(t)\varphi_{u}(t)=\lambda_{u}(t)\,\varphi_{u}(t),\qquad u\in\mathcal{U}(t),$
denote an orthonormal eigenbasis of $M(t)$. Expanding the error in this moving basis gives
$e_{t}=\sum_{u}g_{u}(t)\,\varphi_{u}(t),\qquad g_{u}(t)=\langle e_{t},\varphi_{u}(t)\rangle.$
Differentiating in time yields
$\partial_{t}g_{u}(t)=\langle\partial_{t}e_{t},\,\varphi_{u}(t)\rangle+\big\langle e_{t},\,\partial_{t}\varphi_{u}(t)\big\rangle.$
Using $\dot{e}_{t}=-M(t)e_{t}$, the first term becomes
$\langle-M(t)e_{t},\,\varphi_{u}\rangle=-\lambda_{u}(t)g_{u}(t).$
The second term encodes the rotation of the eigenbasis. From Kato’s perturbation theory (kato2012short; Zwiebach_Adiabatic) for differential self-adjoint operators, the evolution of the eigenfunctions satisfies
$\langle\partial_{t}\varphi_{v}(t),\,\varphi_{u}(t)\rangle=\begin{cases}\displaystyle\frac{\langle\varphi_{u}(t),\,\dot{M}(t)\,\varphi_{v}(t)\rangle}{\lambda_{v}(t)-\lambda_{u}(t)},&v\neq u,\\[10.00002pt] 0,&v=u,\end{cases}$
where $\dot{M}(t)$ is the operator derivative of $M(t)$. Therefore
$\big\langle e_{t},\,\partial_{t}\varphi_{u}(t)\big\rangle=\sum_{v\neq u}g_{v}(t)\,\frac{\langle\varphi_{u}(t),\,\dot{M}(t)\,\varphi_{v}(t)\rangle}{\lambda_{v}(t)-\lambda_{u}(t)}.$
Combining both contributions, we obtain the exact coupled mode ODE:
$\boxed{\partial_{t}g_{u}(t)+\sum_{v\neq u}g_{v}(t)\,\Omega_{v\to u}(t)=-\lambda_{u}(t)\,g_{u}(t),}$
where the coupling coefficients are
$\Omega_{v\to u}(t)=\frac{\langle\varphi_{u}(t),\,\dot{M}(t)\,\varphi_{v}(t)\rangle}{\lambda_{v}(t)-\lambda_{u}(t)}.$
This ODE system is fully rigorous and contains all aspects of the dynamics: local dissipation $\lambda_{u}g_{u}$, nonlocal mode coupling $\sum_{v\neq u}g_{v}\,\Omega_{v\to u}$, and the drift of the feature basis through the time dependence of the eigenfunctions $\varphi_{u}(t)$.
3.2 Logarithmic Spectral Shells and Shell Energies
The exact modewise dynamics derived in Section 3.1 are fully rigorous but too fine-grained for macroscopic analysis. To expose the coarse structure, we group modes into logarithmic spectral shells and track the quadratic error energy carried by each shell.
Fix a ratio $q>1$ and define a logarithmic partition of the positive spectrum:
$\lambda_{\alpha}:=\lambda_{0}\,q^{\alpha},\qquad S_{\alpha}:=\bigl\{\,u:\ \lambda_{u}(t)\in[\lambda_{\alpha},\lambda_{\alpha+1})\,\bigr\},\qquad\alpha\in\mathbb{Z}.$
We also define cumulative shells $S_{\leq\alpha}:=\bigcup_{\gamma\leq\alpha}S_{\gamma}$.
For each shell, we define the shell quadratic energy
$E_{\alpha}(t):=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}(t)^{2},\qquad E_{\leq\alpha}(t):=\sum_{\gamma\leq\alpha}E_{\gamma}(t).$
This choice is canonical: the full function-space loss is $\mathcal{L}(t)=\frac{1}{2}\|e_{t}\|_{2}^{2}=\frac{1}{2}\sum_{u}g_{u}(t)^{2}$, so $\{E_{\alpha}\}$ provides an exact energy bookkeeping across spectral shells.
3.3 Exact Cancellation of Shell-Internal Coupling for Quadratic Energy
Recall the exact mode ODE
$\partial_{t}g_{u}(t)+\sum_{v\neq u}g_{v}(t)\,\Omega_{v\to u}(t)=-\lambda_{u}(t)\,g_{u}(t),\qquad\Omega_{v\to u}(t)=\frac{\langle\varphi_{u}(t),\dot{M}(t)\varphi_{v}(t)\rangle}{\lambda_{v}(t)-\lambda_{u}(t)}.$
Since $\dot{M}(t)$ is self-adjoint and the denominator is antisymmetric, we have the strict antisymmetry
$\Omega_{v\to u}(t)=-\Omega_{u\to v}(t),\qquad u\neq v.$
Proposition 1 (Shell-internal cancellation (quadratic energy)).
For any spectral shell $S_{\alpha}$, the shell-internal coupling does not change the quadratic shell energy:
$\sum_{u\in S_{\alpha}}\sum_{v\in S_{\alpha},\ v\neq u}g_{v}(t)\,g_{u}(t)\,\Omega_{v\to u}(t)=0.$
Proof.
Pair terms $(u,v)$ and $(v,u)$ inside the double sum. Using $\Omega_{v\to u}=-\Omega_{u\to v}$, we get
$g_{v}g_{u}\Omega_{v\to u}+g_{u}g_{v}\Omega_{u\to v}=g_{u}g_{v}\bigl(\Omega_{v\to u}+\Omega_{u\to v}\bigr)=0.$
Summing over all unordered pairs in $S_{\alpha}$ yields the claim. ∎
Remark 1 (Interpretation).
Proposition 1 formalizes the key point: shell-internal coupling is a pure redistribution (rotation) of error among modes within the shell. It can change the individual $g_{u}$’s, but it cannot change $\sum_{u\in S_{\alpha}}g_{u}^{2}$. Therefore, any change of the shell energy $E_{\alpha}(t)$ must come from (i) local dissipation and (ii) energy exchange with other shells.
3.4 Exact Shell Balance Law for Quadratic Energy
Differentiate $E_{\alpha}(t)=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}^{2}$:
$\frac{d}{dt}E_{\alpha}(t)=\sum_{u\in S_{\alpha}}g_{u}(t)\,\partial_{t}g_{u}(t).$
Substituting the mode ODE yields
$\frac{d}{dt}E_{\alpha}(t)$ $=-\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2}-\sum_{u\in S_{\alpha}}\sum_{v\neq u}g_{v}(t)\,g_{u}(t)\,\Omega_{v\to u}(t).$ (1)
Split the coupling sum into $v\in S_{\alpha}$ and $v\notin S_{\alpha}$. By Proposition 1, the $v\in S_{\alpha}$ contribution cancels. Hence we obtain the exact shell energy balance:
$\boxed{\frac{d}{dt}E_{\alpha}(t)=-\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2}-\sum_{\beta\neq\alpha}\ \sum_{\begin{subarray}{c}u\in S_{\alpha}\\ v\in S_{\beta}\end{subarray}}g_{v}(t)\,g_{u}(t)\,\Omega_{v\to u}(t).}$ (2)
The first term is pure shell dissipation. The second term is pure inter-shell exchange.
3.5 Inter-Shell Fluxes and Discrete Conservation Structure
Define the inter-shell quadratic-energy flux from $S_{\beta}$ to $S_{\alpha}$ by
$\mathcal{F}_{\beta\to\alpha}(t):=-\sum_{\begin{subarray}{c}u\in S_{\alpha}\\ v\in S_{\beta}\end{subarray}}g_{v}(t)\,g_{u}(t)\,\Omega_{v\to u}(t).$
Then equation (2) becomes
$\boxed{\frac{d}{dt}E_{\alpha}(t)=-\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2}+\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t).}$ (3)
Moreover, antisymmetry implies a strict action–reaction identity:
$\mathcal{F}_{\beta\to\alpha}(t)=-\mathcal{F}_{\alpha\to\beta}(t).$
Therefore, coupling conserves total quadratic energy across shells:
$\sum_{\alpha}\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t)=0,$
and the only mechanism that decreases $\sum_{\alpha}E_{\alpha}(t)=\frac{1}{2}\|e_{t}\|_{2}^{2}$ is the dissipation term $\sum_{u}\lambda_{u}g_{u}^{2}$.
3.6 From Exact Shell Balance to Renormalized Dynamics
The balance law (3) is an exact conservation-type bookkeeping equation across logarithmic shells, requiring no continuum limit and no PDE interpretation. However, the fluxes $\mathcal{F}_{\beta\to\alpha}(t)$ remain nonlocal and depend on microscopic details of the evolving operator.
In the remainder of the paper, we do not attempt to characterize $\mathcal{F}_{\beta\to\alpha}$ microscopically. Instead, Section 4 introduces a renormalizable shell-dynamics assumption that closes the cumulative effect of inter-shell fluxes at the shell level. All subsequent results rely exclusively on this shell-level interface.
We emphasize that all subsequent results in this paper rely exclusively on this shell-level energy balance and the renormalizable flux interface introduced in Section 4. The continuum PDE description is used only as a convenient approximation when shell resolution permits.
4 Renormalizable Shell Dynamics and Effective Power-Law Transport
Section 3 provides an exact modewise ODE system and an effective spectral PDE description. In this section we introduce an explicitly coarse-grained interface for macroscopic analysis: logarithmic spectral shells, a renormalizable flux bookkeeping condition across shells, and a single effective transport assumption used in the remainder of the paper. We emphasize that this section is not a microscopic characterization of all possible operators $M(t)$; rather, it formulates the minimal coarse-grained structure needed to derive self-similar spectral dynamics and scaling laws.
4.1 Log-shell partition and quadratic shell energies
Fix a ratio $q>1$ and define a logarithmic partition $\{\lambda_{\alpha}\}_{\alpha\in\mathbb{Z}}$ by
$\lambda_{\alpha}:=\lambda_{0}q^{\alpha},\qquad S_{\alpha}:=\{u:\ \lambda_{u}(t)\in[\lambda_{\alpha},\lambda_{\alpha+1})\},\qquad S_{\leq\alpha}:=\bigcup_{\gamma\leq\alpha}S_{\gamma}.$
We work with the quadratic shell energies introduced in Section 3.2:
$E_{\alpha}(t):=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}(t)^{2},\qquad E_{\leq\alpha}(t):=\sum_{\gamma\leq\alpha}E_{\gamma}(t).$
This choice is canonical because $\mathcal{L}(t)=\frac{1}{2}\|e_{t}\|_{2}^{2}=\sum_{\alpha}E_{\alpha}(t)$.
4.2 Exact shell bookkeeping and the role of dissipation
Section 3.4 already established the exact identity
$\frac{d}{dt}E_{\alpha}(t)=-D_{\alpha}(t)+\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t),\qquad D_{\alpha}(t):=\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2},$
where $\mathcal{F}_{\beta\to\alpha}(t)$ is the inter-shell quadratic-energy flux and satisfies $\mathcal{F}_{\beta\to\alpha}(t)=-\mathcal{F}_{\alpha\to\beta}(t)$. In particular, shell-internal coupling does not change $E_{\alpha}$ and the only mechanism that decreases $\sum_{\alpha}E_{\alpha}(t)$ is dissipation $\sum_{\alpha}D_{\alpha}(t)$.
4.3 Definition: renormalizable shell dynamics via cumulative energy flux
The notion of renormalizability adopted here follows the standard usage in statistical physics and field theory. Rather than requiring microscopic locality or exact continuum limits, external interaction effects on a resolution shell are integrated out and absorbed into effective inter-shell fluxes, while subleading corrections become irrelevant at coarse scales. This philosophy underlies Wilsonian renormalization group theory (wilson1983renormalization; kadanoff1966scaling), shell models of turbulence and energy cascades (kolmogorov1995turbulence), and effective hydrodynamic descriptions of nonequilibrium systems (forster1977large; spohn2012large). Our definition formalizes this principle at the level of spectral-shell energy dynamics.
Definition 4.1 (Renormalizable spectral-shell dynamics (energy form)).
The modewise dynamics are said to be (weakly) renormalizable with respect to the log-shell partition if, for sufficiently large $\alpha$, the shell energies admit a closed balance of the form
$\frac{d}{dt}E_{\alpha}(t)=-D_{\alpha}(t)+\mathcal{F}_{\alpha}^{\rm(net)}(t)+\mathcal{R}_{\alpha}(t),$
where the coarse-grained interaction contribution admits a net-flux form
$\mathcal{F}_{\alpha}^{\rm(net)}(t)=J_{\leq\alpha}(t)-J_{\leq\alpha-1}(t),$
and the remainder is negligible compared to the leading dissipation:
$|\mathcal{R}_{\alpha}(t)|\ll D_{\alpha}(t)\qquad\text{as }\alpha\to\infty.$
Here $J_{\leq\alpha}(t)$ is the cumulative quadratic-energy flux across the shell boundary at $\alpha$, defined by the exact microscopic inter-shell fluxes as
$\boxed{J_{\leq\alpha}(t):=\sum_{\beta>\alpha}\ \mathcal{F}_{\beta\to(\leq\alpha)}(t)=\sum_{\beta>\alpha}\ \sum_{\gamma\leq\alpha}\mathcal{F}_{\beta\to\gamma}(t).}$
Moreover, the cumulative flux is integrable at coarse scales in the sense that for each shell $\alpha$ there exists $T_{\alpha}>0$ such that for all $T>T_{\alpha}$,
$\int_{0}^{T}|J_{\leq\alpha}(t)|\,dt\;\leq\;C\int_{0}^{T}D_{\alpha}(t)\,dt.$
Remark 2 (What the cumulative flux represents).
Definition 4.1 does not assume microscopic locality of coupling. All shells $\beta>\alpha$ are included in $J_{\leq\alpha}$. Renormalizability asserts only that after coarse-graining the net inter-shell exchange can be summarized by a controlled boundary flux plus a negligible remainder.
Remark 3 (No direction is assumed).
Renormalizability imposes no restriction on the sign or direction of $J_{\leq\alpha}(t)$. All later results depend only on the existence of a controlled net-flux representation, not on an a priori cascade direction.
4.4 PDE-approximability across shells (energy continuum)
When the shell spacing is sufficiently fine over a resolution range, we may introduce an energy density $\varepsilon(\lambda,t)$ such that
$E_{\alpha}(t)\approx\int_{\lambda_{\alpha}}^{\lambda_{\alpha+1}}\varepsilon(\lambda,t)\,d\lambda,\qquad J_{\leq\alpha}(t)-J_{\leq\alpha-1}(t)\approx\int_{\lambda_{\alpha}}^{\lambda_{\alpha+1}}J(\lambda,t)\,d\lambda,$
This is a local continuum approximation over that range; we do not require a global continuous-spectrum limit. As in nonequilibrium statistical physics forster1977large; spohn2012large, we do not assume the existence of an exact continuum limit; rather, renormalizability refers to the existence of a closed effective description after coarse-graining. We stress again that our continuum equations do not assume the existence of a genuine continuous-spectrum limit. Under this approximation we acquire the expression of the total loss function:
$\mathcal{L}(t)\approx\int_{\lambda_{\min}}^{\lambda_{\max}}\varepsilon(\lambda,t)\,d\lambda.$
They represent a local approximation of the discrete shell-flux difference when shells are sufficiently dense over a dynamically relevant resolution range. Formally, this amounts to approximating the discrete net-flux difference $J_{\leq\alpha}(t)-J_{\leq\alpha-1}(t)$ by a first-order finite-difference representation of $\partial_{\lambda}J$ over a dense but finite shell range. This energy continumm yields, to leading order, the energy transport–dissipation equation
$\partial_{t}\varepsilon(\lambda,t)+\partial_{\lambda}J(\lambda,t)=-2\lambda\,\varepsilon(\lambda,t),$ (4)
on the relevant resolution range. The factor $2$ reflects the fact that dissipation in the mode ODE is $\lambda_{u}g_{u}$, hence quadratic energy dissipates as $\frac{d}{dt}\frac{1}{2}g_{u}^{2}=\lambda_{u}g_{u}^{2}$.
4.5 RMS amplitude notation
For compatibility with earlier GSD/GRSD forms, we also define an RMS amplitude density
$g(\lambda,t):=\sqrt{2\,\varepsilon(\lambda,t)}.$ (5)
Then
$\mathcal{L}(t)\approx\int_{\lambda_{\min}}^{\lambda_{\max}}\varepsilon(\lambda,t)\,d\lambda=\int_{\lambda_{\min}}^{\lambda_{\max}}g(\lambda,t)^{2}\,d\lambda,$
and equation (4) can be viewed as the energy-level backbone behind the GRSD tail forms written in terms of $g$.
4.6 Unified view of NTK and feature learning through spectral transport
The shell-energy formulation provides a clean unification of lazy (NTK-like) training and feature learning through a single macroscopic quantity: the spectral drift velocity $v(\lambda,t)$. Recall the definition
$J(\lambda,t)=v(\lambda,t)\,g(\lambda,t),$
which defines the effective transport velocity $v(\lambda,t)$ in the coarse-grained description.
Lazy / NTK regime ($v\equiv 0$).
If the kernel operator is effectively frozen during training ($\dot{M}(t)\approx 0$), then $v(\lambda,t)\equiv 0$ and no inter-shell transport occurs. The shell dynamics reduce to
$\frac{d}{dt}E_{\alpha}(t)=-D_{\alpha}(t),$
so each shell decays independently under dissipation. No resolution frontier forms, and the dynamics coincide with classical NTK theory.
Feature learning regime ($v\neq 0$).
When the representation evolves ($\dot{M}(t)\neq 0$), the drift velocity becomes active and induces inter-shell energy transport. Energy is redistributed across resolutions before being dissipated, producing a moving resolution frontier and the GRSD tail described in Sections 5–6.
Crucially, this redistribution preserves total quadratic energy and does not increase training loss. Its observable consequences depend on how the transported energy aligns with the evaluation distribution (training vs. test).
Continuum of regimes.
Between these extremes lies a continuum:
  • If $|v(\lambda,t)|\ll\lambda$, dissipation dominates and training is effectively NTK-like.
  • If $|v(\lambda,t)|$ is comparable to or larger than $\lambda$ over a resolution range, transport reshapes the spectrum and feature learning emerges.
Thus, lazy training and feature learning are not distinct dynamical theories, but limiting behaviors of the same renormalized shell-energy dynamics.
Interpretation.
From this perspective, feature learning corresponds to spectral energy transport induced by representation drift, while NTK corresponds to the degenerate zero-transport limit. Both are unified within the same operator-theoretic framework.
4.7 Summary
The key structural takeaway is:
Shell-internal coupling is a rotation that preserves quadratic energy within each shell. Any macroscopic evolution of shell energies is fully captured by (i) dissipation and (ii) inter-shell boundary flux.
All scaling-law consequences in later sections are derived from the renormalized boundary-flux interface (4.1)–(4) in a high-resolution range.
5 Scale-Free GRSD Under Power-Law–Compatible Transport
In this section we analyze the effective GRSD regime arising in an intermediate scale-free training window. Empirically, modern deep models exhibit extended intervals $t_{0}<t<T$ during which macroscopic spectral observables evolve as straight lines in log–log coordinates, indicating the absence of intrinsic scales.
Within such a window, the effective spectral transport must itself be approximately scale-free. While the most general scale-free drift takes the form $v(\lambda,t)=\phi(\lambda/t)$, requiring pure power-law temporal scaling of observables rules out generic $\phi$ and selects the linear scale-free transport
$v(\lambda,t)=c_{0}\,\frac{\lambda}{t},\qquad t_{0}<t<T,$ (6)
up to subleading corrections. In the remainder of this section we take (6) as an effective model assumption and solve the resulting GRSD dynamics explicitly.
Why scale-free observables rigidly constrain the drift.
We emphasize that the power-law form
$v(\lambda,t)=c_{0}\,\frac{\lambda}{t}$
is not introduced as an arbitrary modeling choice. Rather, it is rigidly selected by the requirement that macroscopic spectral observables remain scale-free over an extended training window.
Empirically, in this regime, shell-resolved quantities such as the spectral energy density, cumulative shell energies, and total loss evolve as straight lines in log–log coordinates over multiple decades of time and resolution. This behavior implies not merely scale invariance at a fixed time, but a joint scale covariance between the spectral coordinate $\lambda$ and the training time $t$.
At the level of the coarse-grained shell dynamics, this covariance severely restricts the admissible form of the effective transport velocity. While the most general scale-free drift compatible with dimensional analysis can be written as $v(\lambda,t)=\phi(\lambda/t)$, generic choices of $\phi$ induce nonlinear reparameterizations of the spectral coordinate and therefore produce curvature in log–log observables. Requiring that these observables remain asymptotically affine in log–log space selects the linear form $\phi(u)=c_{0}u$, yielding $v(\lambda,t)=c_{0}\lambda/t$ up to subleading corrections.
Lagrangian spectral coordinates as the renormalized frame.
The emergence of the Lagrangian spectral coordinate $u$ should not be viewed as a technical device introduced solely to solve the transport equation. Instead, it represents the natural renormalized variable induced by the shell-level flux interface.
At the discrete shell level, renormalizability asserts that the cumulative effect of inter-shell interactions can be summarized by a controlled boundary flux. In the continuum approximation, this structure manifests as a transport term that can be absorbed into a time-dependent reparameterization of the spectral coordinate. The characteristic flow (5) implements precisely this absorption.
In the co-moving coordinate $u$, the dynamics reduce to pure dissipation. All nontrivial macroscopic structure—including scaling laws and apparent emergence— is therefore encoded in how this co-moving frame maps back to stationary spectral coordinates. The simplicity of the resulting evolution is not accidental, but a direct consequence of the renormalizable shell-dynamics assumption.
5.1 Characteristic flow and scale-free reparameterization
Starting from the leading-order GRSD equation
$\partial_{t}\varepsilon(\lambda,t)+\partial_{\lambda}\!\big(v(\lambda,t)\varepsilon(\lambda,t)\big)=-2\lambda\,\varepsilon(\lambda,t),$ (7)
with drift (6), the characteristic curves satisfy
$\frac{d\lambda}{dt}=c_{0}\,\frac{\lambda}{t}.$ (8)
Integrating yields the exact scale-free flow
$\lambda(t;u)=u\Big(\frac{t}{t_{0}}\Big)^{c_{0}},\qquad u=\lambda\Big(\frac{t}{t_{0}}\Big)^{-c_{0}},$ (9)
where $u$ defines a stationary (Lagrangian) spectral coordinate. This flow is strictly power-law and affine in log–log coordinates.
5.2 Energy density along characteristics
Along a characteristic $\lambda(t;u)$, the energy density evolves as
$\frac{d}{dt}\log\varepsilon(\lambda(t),t)=-2\lambda(t).$ (10)
Substituting (9) and integrating gives
$\varepsilon(u,t)=\varepsilon_{0}(u)\,\exp\!\Big[-\frac{2u}{1+c_{0}}\Big(\Big(\frac{t}{t_{0}}\Big)^{1+c_{0}}-1\Big)\Big],\qquad c_{0}\neq-1.$ (11)
which is aligned with the spectral evolution conjecture proposed in zhang2025generalized:
$f^{*}-f_{\lambda}(t)=w_{\lambda}\exp(g(\lambda,t))$ (12)
where $g(\lambda,t)\propto\lambda^{a(\beta)}t^{b(\beta)}$ and thus $w_{\lambda}^{2}$ correspond to the starting error on mode with eigenvalue of $\lambda$. This conjecture is helpful for understanding the scaling law for both loss evolution and model compression. It further predicted the existence of learning frontier and defines model density, which is observed to evolve according to Xiao2024DensingLaw.
Rewriting (11) in Eulerian coordinates using $u=\lambda(t/t_{0})^{-c_{0}}$ yields
$\boxed{\varepsilon(\lambda,t)=\varepsilon_{0}\!\big(\lambda\,t^{-c_{0}}\big)\,\exp\!\Big(-\frac{2\lambda}{1+c_{0}}\,t\Big)}.$ (13)
The exponential factor represents physical dissipation, while the prefactor encodes pure scale-free transport. The competition between transport and dissipation is governed entirely by the combination $1+c_{0}$. This expression constitutes the leading-order GRSD solution in the scale-free window.
The total quadratic loss is
$\mathcal{L}(t)=\int_{0}^{\infty}\varepsilon(\lambda,t)\,d\lambda.$
Substituting (13) and performing a standard moving-front estimate yields power-law decay
$\mathcal{L}(t)\sim t^{-(1+c_{0})(1-\alpha)},$
for initial spectra $\varepsilon_{0}(u)\sim u^{-\alpha}$.
Transport–dissipation competition and spectral front formation.
The explicit solution reveals a universal structural feature: the dynamics are governed by a competition between scale-free transport and scale-dependent dissipation. Transport alone redistributes error mass across resolutions without loss, while dissipation selectively suppresses high-$\lambda$ components.
Their interplay generates a moving spectral front separating a learned region from an unresolved tail. Importantly, this front is not an artifact of a particular initialization, architecture, or optimizer choice. It is a robust consequence of renormalizable spectral transport under power-law–compatible drift.
In the next section, we show that when training and evaluation are performed on finite samples drawn from the same underlying distribution, this moving front interacts with unavoidable high-resolution spectral mismatch. The resulting transient amplification in certain spectral bands provides a natural and entirely in-domain explanation of the double-descent phenomenon.
6 Double Descent from In-Domain Finite-Sample Spectral Mismatch
We now explain the origin of the double-descent phenomenon from the perspective of renormalizable spectral-shell dynamics. Crucially, the mechanism described here is entirely in-domain: training and test data are drawn independently from the same underlying data-generating distribution. The non-monotonicity of the test loss arises not from distribution shift, but from unavoidable finite-sample spectral mismatch concentrated in high-resolution modes.
6.1 Training dynamics and monotone training loss.
Throughout training, the error evolves according to
$\dot{e}_{t}=-M(t)e_{t},\qquad M(t)=J_{\theta(t)}J_{\theta(t)}^{\!*}\succeq 0.$
Measured in the training metric, the quadratic loss
$\mathcal{L}_{\rm tr}(t):=\frac{1}{2}\|e_{t}\|_{L^{2}(p)}^{2}=\sum_{\alpha}E_{\alpha}(t),\qquad E_{\alpha}(t):=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}(t)^{2},$
is strictly non-increasing:
$\frac{d}{dt}\mathcal{L}_{\rm tr}(t)=-\langle e_{t},M(t)e_{t}\rangle\leq 0.$
This monotonicity is exact and does not rely on any coarse-graining, continuum approximation, or weak-coupling assumption.
6.2 In-domain train/test mismatch as a finite-sample effect.
Let $P_{\rm tr}$ and $P_{\rm te}$ denote the empirical quadratic forms induced by the training and test samples, respectively. Both are unbiased estimators of the same population operator associated with the underlying distribution, but differ due to finite sampling. Define the mismatch operator
$\Delta:=P_{\rm te}-P_{\rm tr}.$
The test loss can be written as
$\mathcal{L}_{\rm te}(t)=\frac{1}{2}\langle e_{t},P_{\rm te}e_{t}\rangle=\mathcal{L}_{\rm tr}(t)+\frac{1}{2}\langle e_{t},\Delta e_{t}\rangle.$
The second term captures purely finite-sample effects and vanishes only in the infinite-data limit.
For continuous data distributions or infinite-rank kernel operators, the population spectrum exhibits vanishing eigenvalue gaps in the spectral tail. Classical results on empirical covariance and integral operators show that, in this regime, the associated empirical eigenspaces exhibit $O(1)$ fluctuations under finite sampling (koltchinskii2017normal; rosasco2010learning). Consequently, even under in-domain sampling, $\Delta$ is generically supported on high-resolution (spectral-tail) components and is not co-diagonalizable with the training-induced operator $M(t)$.
Double descent as a generic finite-sample effect.
We stress that the non-monotonic behavior described here does not rely on distribution shift, adversarial sampling, or overparameterization per se. It arises generically whenever the population operator exhibits a slowly decaying spectral tail and learning proceeds through a transport-dominated regime.
From this perspective, double descent is not a pathological deviation from monotone learning, but a predictable finite-sample manifestation of renormalizable spectral transport. As model capacity or dataset size increases, the location of the mismatch-dominated spectral region shifts, modifying the timing and magnitude of the transient amplification without eliminating the underlying mechanism.
6.3 Transport-induced amplitude reweighting
Starting from the GRSD solution (13), the spectral error density admits the exact factorization
$\varepsilon(\lambda,t)=A(\lambda,t)\,\exp\!\Big(-\frac{2\lambda}{1+c_{0}}\,t\Big),\qquad A(\lambda,t):=\varepsilon_{0}(\lambda t^{-c_{0}}).$ (14)
The exponential factor represents pure dissipation and is strictly decreasing in time. All non-monotonic behavior at fixed $\lambda$ is therefore encoded in the amplitude function $A(\lambda,t)$, which arises solely from scale-free transport.
Intuitive picture.
Intuitively, scale-free transport continuously transfers error mass from higher-$\lambda$ modes, which are learned and dissipated rapidly, toward lower-$\lambda$ modes. As a result, even as high-resolution errors decay, part of this error is temporarily deposited onto lower-resolution components, raising their spectral error density.
At the same time, dissipation at scale $\lambda$ is proportional to $\lambda\,\varepsilon(\lambda,t)$, and is therefore weaker at small $\lambda$. The mismatch between fast transport from high $\lambda$ and delayed dissipation at low $\lambda$ naturally produces a transient amplification before eventual decay.
Differentiating the amplitude function at fixed $\lambda$ yields
$\partial_{t}A(\lambda,t)=-\,c_{0}\,\lambda\,t^{-c_{0}-1}\,\varepsilon_{0}^{\prime}(\lambda t^{-c_{0}}).$ (15)
Hence, whenever the initial spectrum is decreasing, $\varepsilon_{0}^{\prime}(u)<0$, and $c_{0}>0$, the amplitude grows in time:
$\partial_{t}A(\lambda,t)>0.$
Consequently, the spectral error density $\varepsilon(\lambda,t)$ can increase transiently at fixed $\lambda$, despite overall dissipation and monotone decay of the training loss.
6.4 Double descent from transport-induced amplification in the mismatch window
We now formalize how transport-induced amplitude growth leads to double descent when combined with finite-sample spectral mismatch. The key point is that, although the training loss is strictly monotone, the test loss probes a different quadratic form that selectively weights high-resolution spectral components.
Spectral localization of finite-sample mismatch.
Recall that the test loss can be written as
$L_{\mathrm{te}}(t)=L_{\mathrm{tr}}(t)+\frac{1}{2}\langle e_{t},\Delta e_{t}\rangle,\qquad\Delta:=P_{\mathrm{te}}-P_{\mathrm{tr}},$
where $\Delta$ is the mismatch operator. Classical results on empirical covariance and integral operators imply that, for continuous data distributions and infinite-rank population operators, $\Delta$ is generically supported on high-resolution modes, corresponding to small eigenvalues $\lambda$ of the population operator.
At the coarse-grained level, this implies that there exists a resolution interval
$\lambda\in[\lambda_{-},\lambda_{+}],$
which we refer to as the mismatch window, such that the dominant contribution to $\langle e_{t},\Delta e_{t}\rangle$ comes from spectral components in this range. Outside this window, either the modes are well-aligned between train and test or their contribution is suppressed by dissipation.
Mismatch-weighted test loss and non-monotonicity.
The mismatch contribution to the test loss can be approximated, to leading order, by restricting to the mismatch window:
$\langle e_{t},\Delta e_{t}\rangle\;\approx\;\int_{\lambda_{-}}^{\lambda_{+}}w(\lambda)\,\varepsilon(\lambda,t)\,d\lambda,$
where $w(\lambda)\geq 0$ denotes the coarse-grained spectral weight induced by the finite-sample mismatch operator $\Delta=P_{\mathrm{te}}-P_{\mathrm{tr}}$ in the instantaneous spectral basis of $M(t)$, so that $w(\lambda)\,d\lambda$ represents the leading-order contribution of modes with eigenvalues in $\lambda,\lambda+d\lambda]$ to the quadratic form $\langle e_{t},\Delta e_{t}\rangle$. Formally, $w(\lambda)$ is defined by coarse-graining the diagonal spectral density of the mismatch operator $\Delta$ in the instantaneous eigenbasis of $M(t)$, i.e.
$w(\lambda)\;:=\;\sum_{u:\,\lambda_{u}(t)\in[\lambda,\lambda+d\lambda)}\langle\phi_{u}(t),\Delta\,\phi_{u}(t)\rangle\,/\,d\lambda,$
so that $\langle e_{t},\Delta e_{t}\rangle\approx\int w(\lambda)\,\varepsilon(\lambda,t)\,d\lambda$ to leading order.
Substituting the factorized form of $\varepsilon$ gives
$\int_{\lambda_{-}}^{\lambda_{+}}w(\lambda)\,A(\lambda,t)\,\exp\!\left(-\frac{2\lambda}{1+c_{0}}\,t\right)\,d\lambda.$
If the mismatch window $\lambda_{-},\lambda_{+}]$ intersects the region of spectral coordinates for which $\partial_{t}A(\lambda,t)>0$, then the integrand exhibits a competition between transport-induced amplification and dissipation. At early times, the growth of $A(\lambda,t)$ dominates, leading to an increase of the mismatch contribution. At later times, exponential dissipation takes over, forcing eventual decay.
Sufficient condition for double descent.
We therefore obtain the following sufficient condition: if there exists a time interval and a non-negligible spectral region
$\lambda\in[\lambda_{-},\lambda_{+}]\quad\text{such that}\quad\partial_{t}A(\lambda,t)>0,$
then the test loss $L_{\mathrm{te}}(t)$ is necessarily non-monotone, even though the training loss $L_{\mathrm{tr}}(t)$ decreases monotonically. This non-monotonicity manifests as the classical double-descent curve.
Importantly, this mechanism is entirely in-domain. It does not rely on distribution shift or adversarial effects, but follows generically from the interaction between renormalizable spectral transport and finite-sample spectral mismatch.
7 Related Work
Neural scaling laws.
Empirical studies have established that neural networks display remarkably regular power-law relationships between compute, model size, dataset size, and achievable loss (kaplan2020scaling; henighan2020scaling; hestness2017deep; hoffmann2022training; hernandez2021scaling). Recent work has investigated both the emergence of such scaling behavior and its theoretical underpinnings, including dynamical models of loss evolution (bordelon2024dynamical) and the role of data pruning (sorscher2022beyond). Precision scaling laws have also been explored in the context of architectural and quantization constraints (kumar2024scaling).
Feature learning and spectral bias.
A major line of work studies how neural networks acquire hierarchical representations and exhibit spectral preference (rahaman2019spectral; bordelon2023feature). Kernel-based analyses of learning dynamics (bietti2021inductive; canatar2022kernel) and infinite-width approximations (jacot2018neural; lee2019wide; yang2021tensor) have contributed significantly to understanding the transition between lazy training and representation learning. Recent studies have also revealed consistency of learned features across widths (vyas2023feature) and the spectral evolution of networks (wang2023spectralevolution).
Optimization dynamics and stability.
The geometry of optimization landscapes, the effect of batch size, and training stability have been examined extensively (keskar2017large; ghorbani2019investigation). Work on pruning and compression (rosenfeld2021predictability; han2015deep; blalock2020state; lecun1990optimal; molchanov2017variational; lee2019snip; wang2020picking; han2016deep; nagel2021up; frantar2022gptq; dettmers2022llm) has illuminated how spectral structure interacts with parameter sparsity and low-precision computation. Theoretical connections between spectral evolution, implicit bias, and neural dynamics continue to be an active area of research (domine2024lazy; zhang2025generalized).
Perturbation theory and adiabatic analysis.
Our spectral formulation draws on classical operator perturbation theory (kato2012short) and its analogues in quantum adiabatic evolution (Zwiebach_Adiabatic). These tools formalize how eigenfunctions and eigenvalues evolve under smooth or weakly coupled updates, providing a principled foundation for the drift–dissipation dynamics developed in this work.
8 Conclusion
This work proposed a spectral–shell framework for understanding neural scaling laws, feature learning, and double-descent phenomena directly from the operator-level dynamics induced by gradient descent. Starting from the exact function-space evolution $\dot{e}_{t}=-M(t)e_{t}$, we derived a rigorous modewise formulation and showed that, upon logarithmic coarse-graining, shell-internal interactions cancel identically at the level of quadratic error energy. As a result, the macroscopic evolution of error is governed entirely by dissipation and inter-shell energy exchange.
The central modeling ingredient of this paper is a renormalizable spectral-shell dynamics assumption: after coarse-graining, the cumulative effect of microscopic interactions can be summarized by a controlled net energy flux across shell boundaries. Under an effective power-law form of this renormalized shell-level flux, the shell dynamics admit a self-similar high-resolution tail with a moving resolution frontier. This structure yields explicit scaling-law behavior for the total loss and provides a unified explanation of neural scaling laws and double descent.
A key conceptual outcome of this framework is a unified view of lazy (NTK-like) training and feature learning. When inter-shell transport is negligible, shell energies decay independently under dissipation, recovering classical kernel dynamics. When transport is active, spectral energy is redistributed across resolutions before being dissipated, inducing representation learning. Both regimes—and the continuum between them—are governed by the same shell-energy bookkeeping, differing only in the effective shell-flux interface.
Importantly, the analysis does not assume the existence of a continuous spectrum or a vanishing shell spacing. All continuum partial differential equations appearing in the paper serve only as convenient local approximations of discrete shell-energy dynamics over resolution ranges where shells are sufficiently dense. All scaling-law conclusions can be equivalently interpreted at the discrete shell level.
Future directions.
Several important questions remain open:
  1. 1.
    Multi-task and multi-distribution learning. The present analysis focuses on a single task and a single data distribution. Extending spectral-shell dynamics to multi-task settings—where different task operators may not share eigenbases and may interact through shared representation drift—could reveal new mechanisms of transfer, interference, and modular generalization.
  2. 2.
    Origin of renormalizability and power-law shell transport. In this work, the renormalizable shell-flux interface and its effective power-law form are treated as macroscopic assumptions. A natural next step is to understand their qualitative and quantitative origins from operator-level structure. Possible mechanisms include locality of spectral coupling, effective one-directional energy transfer across resolution shells, and the suppression of long-range interactions by stochastic optimization noise. Clarifying when and why such mechanisms produce renormalizable and power-law shell dynamics would substantially deepen the theoretical foundations of the framework.
  3. 3.
    Modeling optimizers and learning-rate schedules. The present formulation is developed for gradient-based optimization with an effective time reparameterization. Momentum, adaptive methods, and second-order or approximate natural-gradient schemes can modify both dissipation and transport. Developing spectral-shell descriptions of these optimizers—and of common learning-rate schedules—may explain optimizer-dependent variations in scaling behavior.
  4. 4.
    Connecting $J_{\theta}$ to network architecture. The Jacobian operator $J_{\theta}$ mediates how parameters generate spectral structure in function space. Understanding how architectural features such as depth, width, and parameterization shape the induced shell dynamics could connect model scaling, representational capacity, and achievable loss within a unified operator-theoretic framework.
Overall, the spectral-shell perspective developed here isolates a minimal and robust macroscopic structure underlying neural training dynamics. By separating exact shell-level conservation laws from coarse-grained flux assumptions, it provides a flexible foundation for understanding scaling phenomena while leaving room for future work on their microscopic origins.
Appendix A Exact Shell-Energy Bookkeeping from Modewise Dynamics
This appendix derives the exact shell-level energy balance underlying the main text. Its purpose is purely structural: to show how the quadratic shell energies used throughout the paper arise rigorously from the exact modewise dynamics. No weak-coupling, locality, or continuum assumptions are imposed here. All scaling-law results rely only on the renormalizable shell-dynamics assumption stated in Definition 4.1.
A.1 Modewise dynamics and antisymmetric coupling
Recall from Section 3.1 that, in the instantaneous eigenbasis of $M(t)$, the error amplitudes satisfy the exact system
$\partial_{t}g_{u}(t)+\sum_{v\neq u}g_{v}(t)\,\Omega_{v\to u}(t)=-\lambda_{u}(t)\,g_{u}(t),$ (16)
with coupling coefficients
$\Omega_{v\to u}(t)=\frac{\langle\varphi_{u}(t),\dot{M}(t)\varphi_{v}(t)\rangle}{\lambda_{v}(t)-\lambda_{u}(t)},\qquad u\neq v.$
Since $\dot{M}(t)$ is self-adjoint, these coefficients satisfy the strict antisymmetry
$\Omega_{v\to u}(t)=-\Omega_{u\to v}(t),\qquad u\neq v.$ (17)
A.2 Logarithmic shells and quadratic energies
Fix a logarithmic partition of the spectrum as in Section 3.2:
$S_{\alpha}:=\{u:\ \lambda_{u}(t)\in[\lambda_{\alpha},\lambda_{\alpha+1})\},\qquad\lambda_{\alpha}=\lambda_{0}q^{\alpha},\quad q>1.$
For each shell, define the quadratic shell energy
$E_{\alpha}(t):=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}(t)^{2}.$ (18)
This choice is canonical, since the total training loss is exactly
$\mathcal{L}(t)=\frac{1}{2}\|e_{t}\|_{2}^{2}=\sum_{\alpha}E_{\alpha}(t).$
A.3 Exact cancellation of shell-internal coupling
Differentiate (18):
$\frac{d}{dt}E_{\alpha}(t)=\sum_{u\in S_{\alpha}}g_{u}(t)\,\partial_{t}g_{u}(t).$
Substituting (16) yields
$\frac{d}{dt}E_{\alpha}(t)=-\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2}-\sum_{u\in S_{\alpha}}\sum_{v\neq u}g_{u}(t)g_{v}(t)\,\Omega_{v\to u}(t).$ (19)
Split the coupling sum into contributions with $v\in S_{\alpha}$ and $v\notin S_{\alpha}$. By antisymmetry (17), the shell-internal contribution vanishes identically:
$\sum_{u\in S_{\alpha}}\sum_{\begin{subarray}{c}v\in S_{\alpha}\\ v\neq u\end{subarray}}g_{u}g_{v}\,\Omega_{v\to u}=0.$ (20)
Thus shell-internal coupling acts as a pure rotation of amplitudes and does not change the quadratic shell energy.
A.4 Exact shell-energy balance
Only dissipation and inter-shell exchange remain. Define the inter-shell quadratic-energy flux
$\mathcal{F}_{\beta\to\alpha}(t):=-\sum_{\begin{subarray}{c}u\in S_{\alpha}\\ v\in S_{\beta}\end{subarray}}g_{u}(t)g_{v}(t)\,\Omega_{v\to u}(t).$
Then (19) reduces to the exact identity
$\boxed{\frac{d}{dt}E_{\alpha}(t)=-\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2}+\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t).}$ (21)
Moreover, antisymmetry implies the action–reaction property
$\mathcal{F}_{\beta\to\alpha}(t)=-\mathcal{F}_{\alpha\to\beta}(t),$
so that inter-shell coupling conserves total quadratic energy:
$\sum_{\alpha}\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t)=0.$
The only mechanism that decreases $\sum_{\alpha}E_{\alpha}(t)$ is dissipation.
A.5 Relation to renormalizable shell dynamics
Equation (21) is an exact bookkeeping identity. It involves no approximation, no continuum limit, and no assumption on the structure of the coupling coefficients beyond antisymmetry.
The renormalizable shell-dynamics assumption introduced in Definition 4.1 is a macroscopic closure of this identity: it postulates that, after coarse-graining, the cumulative effect of the inter-shell fluxes can be summarized by a controlled net boundary flux plus a negligible remainder. All scaling-law results in the main text are derived from this shell-level closure and do not rely on any further microscopic characterization.
Appendix B Subcritical Drift Regimes ($b\leq 1$)
In the main text, we focus on the supercritical regime $b>1$, under which the spectral drift induces a progressive resolution cascade and gives rise to well-defined data scaling laws. For completeness, we briefly discuss the remaining cases $b\leq 1$ here, which correspond to qualitatively different learning dynamics.
Critical regime ($b=1$).
When $b=1$, the characteristic equation
$\frac{d\lambda}{dt}=-c(t)\lambda$
admits an exponential solution $\lambda(t)=\lambda_{0}e^{-\tau(t)}$, where $\tau(t)=\int_{0}^{t}c(s)\,ds$. In this regime, all spectral modes decay at the same relative rate, and no resolution hierarchy is formed. This behavior corresponds to the well-known lazy or NTK regime, in which the induced kernel remains effectively fixed and feature learning is absent. While learning still occurs, it does not exhibit progressive resolution unlocking or nontrivial scaling with data.
Subcritical regime ($0<b<1$).
For $0<b<1$, the characteristic flow
$\lambda(t)^{1-b}=\lambda_{0}^{1-b}-(1-b)\tau(t)$
reaches the spectral boundary $\lambda=0$ in finite effective time. In practical settings, this boundary is replaced by a resolution floor determined by finite model capacity, numerical precision, or noise. As a result, the spectral mass is rapidly transported toward this floor, after which no further refinement of resolution is possible.
Importantly, learning is not absent in this regime. Rather, the model converges quickly to a fixed representational limit, beyond which increasing the amount of data only accelerates convergence toward the same solution. In particular, no asymptotic scaling law with respect to data size can be sustained. In this sense, the learning dynamics in the subcritical regime are not scalable.
Implications.
The regimes $b\leq 1$ thus correspond to either non-hierarchical (lazy) or non-scalable learning dynamics. Since modern deep learning systems of interest empirically exhibit strong scalability with data, we restrict our main analysis to the supercritical regime $b>1$, and defer the discussion of subcritical cases to this appendix.
Appendix C The polynomial bound of effective time $\tau(t)$
Theorem 2 (Polynomial Growth of the Effective Time under Standard Optimization Assumptions).
Let the drift strength $c(t)$ of the spectral velocity $v(\lambda,t)\sim-c(t)\lambda^{b}$ be determined by the gradient statistics of a neural network trained with a stable optimizer (SGD, AdamW, momentum SGD) under bounded gradient norms and non-exploding learning-rate schedules. Then the effective time
$\tau(t)=\int_{0}^{t}c(s)\,ds$
satisfies
$\tau(t)=t^{\alpha}\,L(t),\qquad\alpha>0,$
where $L(t)$ is a slowly varying function in the sense of Karamata:
$\lim_{t\to\infty}\frac{L(kt)}{L(t)}=1\qquad\forall k>0.$
In particular, $\tau(t)$ is asymptotically polynomial up to a slowly varying correction.
Proof.
The drift amplitude $c(t)$ is determined by
$v(\lambda,t)=-c(t)\lambda^{b},\qquad c(t)=\langle\phi_{\lambda}(t),\dot{M}(t)\,\phi_{\lambda}(t)\rangle\,\lambda^{-b}.$
The operator derivative $\dot{M}(t)$ equals the expected gradient outer product
$\dot{M}(t)=\mathbb{E}\bigl[\nabla_{\theta}f(x;\theta(t))\,\nabla_{\theta}f(x;\theta(t))^{\top}\bigr]\cdot\eta(t),$
where $\eta(t)$ is the learning rate.
Under standard stability conditions of deep learning:
  • bounded gradient norms: $\|\nabla_{\theta}f(x;\theta(t))\|\leq C$,
  • non-exploding learning rates,
  • Lipschitz continuity of the gradient field,
  • noise-controlled SGD (bounded variance),
the quantity $c(t)$ is positive and satisfies
$c_{\min}\,t^{\alpha-1}\leq c(t)\leq c_{\max}\,t^{\alpha-1}L(t),$
for some $\alpha>0$ and a slowly varying function $L$. This follows from classical results on stochastic approximation and ODE–SDE averaging, which imply that gradient magnitudes evolve according to polynomially bounded envelopes and admit slowly varying corrections through learning-rate modulation and noise.
Integrating $c(t)$, we obtain
$\tau(t)=\int_{0}^{t}c(s)\,ds=t^{\alpha}\,L(t),$
where the same $L(t)$ (up to asymptotic equivalence) is slowly varying because integration preserves slow variation.
Thus the effective time is asymptotically a polynomial multiplied by a slowly varying factor. ∎